### v1

# DISC: Dynamic Decomposition Improves LLM Inference Scaling

Abstract
Inference scaling methods often rely on decomposing problems into steps (or groups of tokens), followed by sampling and selecting the best next steps. However, these steps and their sizes are often predetermined or manually designed based on domain knowledge. We propose dynamic decomposition, a method that adaptively and automatically fractions solution and reasoning traces into manageable steps during inference. By more effectively allocating compute – particularly through subdividing challenging steps and prioritizing their sampling – dynamic decomposition significantly improves inference efficiency. Experiments on benchmarks such as APPS, MATH, and LiveCodeBench demonstrate that dynamic decomposition outperforms static approaches, including token-level, sentence-level, and single-step decompositions. These findings highlight the potential of dynamic decomposition to improve a wide range of inference scaling techniques.
Machine Learning, LLMs, Inference scaling, test-time compute, planning, reasoning, decomposition
1 Introduction
Scaling inference efficiency remains a fundamental challenge for large language models (LLMs). Many existing approaches improve inference by decomposing problems into smaller steps and systematically exploring different solutions (Feng et al., 2023; Zeng et al., 2024; Wu et al., 2024; Nori et al., 2024; Snell et al., 2024; Brown et al., 2024; Gandhi et al., 2024; Lee et al., 2025; Light et al., 2024a; Wang et al., 2025).
Some decomposition methods rely on domain-specific heuristics and hand-crafted rules (Yao et al., 2024; Zelikman et al., 2023; Zhou et al., 2022). However, manually partitioning problems or designing task-specific heuristics is costly and lacks generalization. Moreover, identifying critical steps for an LLM can be non-trivial for humans. As shown in Sec. 3.5, LLMs may assign importance to seemingly trivial words (e.g., therefore or which), which, while counterintuitive to humans, play a crucial role in autoregressive generation (Lin et al., 2025). Other approaches employ fixed, uniform step sizes, such as token- or sentence-level decomposition (Feng et al., 2023; Guo et al., 2025). All these methods rely on static decomposition strategies, where step sizes are predefined or determined via heuristics. Such rigidity wastes compute on steps that are easy for the LLM (but potentially difficult for humans) while undersampling more challenging steps.
To overcome these limitations, we propose DISC (Dynamic decomposition Improves Scaling Compute), a recursive inference algorithm that dynamically partitions solution steps based on difficulty. Unlike prior methods, DISC adapts decomposition granularity during inference based on both the available budget and problem complexity, ensuring finer granularity for more difficult steps. By leveraging the autoregressive nature of LLMs, DISC efficiently locates difficult steps through binary partitioning, focusing compute on challenging regions rather than wasting resources on trivial steps. DISC is generalizable and requires no human supervision, domain-specific heuristics, prompt engineering, or process annotations, making it widely applicable across tasks.
Our main contributions are:
  • We introduce DISC, a method for recursive partitioning and decomposing solutions during inference without human supervision, domain-specific heuristics, or process reward models.
  • We demonstrate how DISC integrates decomposition with inference-time search, allocating compute to high-impact, difficult steps.
  • We show that DISC improves inference scaling in terms of both sample efficiency and token efficiency.
  • We provide insights into how LLMs reason and plan by identifying critical steps in their generation process.
2 Preliminaries
Figure 1: Comparison of different automatic decomposition methods based on step size determination.
2.1 Problem Setting
We consider a reasoning and code generation setting where a dataset $\mathcal{X}=\{\boldsymbol{x}^{(i)}\}_{i=1}^{N}$ consists of problem prompts $\boldsymbol{x}$, and a reward model $R:\mathcal{X}\cdot\mathcal{Y}\rightarrow[0,1]$ evaluates generated solutions $\boldsymbol{y}\in\mathcal{Y}$. This includes program synthesis, where correctness is verified using ground-truth tests (Chen et al., 2021; Austin et al., 2021), and mathematical reasoning, where solutions are validated numerically (Hendrycks et al., 2021a; Cobbe et al., 2021). The reward model can be a ground-truth verifier, a trained heuristic (Zhang et al., 2024), self-consistency (Wang et al., 2023a), or an LLM-as-a-judge (Zheng et al., 2023). Since our focus is on step decomposition rather than verification, we use the ground-truth reward model where available. We assume access to a pretrained language model $\pi$ that generates text autoregressively. A generated response $\boldsymbol{y}$ consists of both the final solution and the reasoning chain leading to it, and can be represented as a sequence of tokens $\boldsymbol{y}=(y_{0},...,y_{L_{\boldsymbol{y}}})$. Additionally, solutions can be partitioned into solution steps $\boldsymbol{y}=(\boldsymbol{y}_{0},...,\boldsymbol{y}_{K})$, where each step $\boldsymbol{y}_{i}$ is a contiguous string of tokens. A partial solution up to step $k$ is defined as $\boldsymbol{y}_{1...k}:=\boldsymbol{y}_{1}\cdot\boldsymbol{y}_{2}\cdot...\cdot\boldsymbol{y}_{k}$, and its rollout or completion, denoted $\boldsymbol{y}_{1...k+}$, is the continuation generated by $\pi$ until an end-of-sequence token (EOS). The size of a solution step, $|\boldsymbol{y}_{i}|$, refers to its length in tokens or characters.
2.2 Prior Automatic Decomposition Methods
Single-step generation. In a single-step generation, the entire solution is generated in one pass from the prompt to the EOS token, treating it as a single action. This approach underlies the widely used inference scaling method best of n (BoN) (Cobbe et al., 2021; Lightman et al., 2023; Snell et al., 2024; Liang et al., 2024), where $n$ complete solutions are sampled, and the highest-scoring one is selected. Single-step generation also plays a role in alignment and fine-tuning methods such as DPO (Rafailov et al., 2024) and RLOO (Ahmadian et al., 2024).
Token-level decomposition. At the opposite end of the spectrum, token-level decomposition treats each atomic token as an individual step. While this approach dramatically increases search complexity, it enables fine-grained search that can yield higher performance gains given sufficient compute (Feng et al., 2023).
Newline and sentence-level decomposition. A commonly used decomposition method segments LLM generations into sentences or lines based on delimiters such as periods or newlines (Hao et al., 2023; Feng et al., 2023; Yao et al., 2024). Typically, each newline corresponds to a new paragraph, equation, or line of code, which often encapsulates a distinct reasoning step.
Problem: Automatic and scalable decomposition Existing decomposition methods are task-specific, manual, and static, limiting their adaptability and scalability.
2.3 Step Sampling
Inference-time scaling methods must balance exploration at the current step with exploration of future steps. We implement a simple dynamic sampling process, referred to as negative binomial sampling, where we continue sampling completions until the sum of their rewards exceeds a predefined threshold $\sigma$. More formally, the number of samples $M$ drawn from a partial solution $\boldsymbol{y}_{1...k}$ is the smallest integer satisfying $\sum_{i=1}^{M}R(\boldsymbol{x}\cdot\boldsymbol{y}_{1...k+}^{(i)})\geq\sigma$.
Where $\boldsymbol{y}_{1...k+}^{(i)}$ represents the $i$-th sampled completion from the partial solution. This process ensures efficient allocation of compute by dynamically adjusting the number of samples per step. It achieves this by either: (a) continuing to sample until a sufficiently high-reward completion is found, or (b) stopping early when additional sampling is unlikely to yield significant improvements, thus redirecting the compute to future steps. The stopping criterion is governed by $\sigma$: when accumulated reward from completions surpasses $\sigma$, the method assumes further sampling is unnecessary. For a fair comparison, we apply this sampling method uniformly across all decomposition methods in our experiments.
2.4 Inference Scaling Methods and Decomposition
Since our study focuses on decomposition rather than search, we primarily use greedy step search as the search method. In greedy step search, multiple candidate steps are sampled at each iteration, but only the highest-scoring step is retained, while the rest are discarded. The process then repeats, conditioning future steps on the best step found so far. We also perform ablation studies comparing Monte Carlo Tree Search (MCTS)  (Feng et al., 2023; Light et al., 2024b) and beam search  (Xie et al., 2024), two commonly used inference scaling methods. These comparisons, presented in Sec. 4.2, highlight how different search strategies interact with decomposition. Additional details on MCTS and beam search are provided in App. E.
3 Methodology
Figure 2: DISC recursively partitions the most challenging step – i.e., the one with the lowest priority metric $h$ – to progressively refine and find the optimal solution.
3.1 DISC Algorithm
The DISC algorithm employs recursive binary decomposition to iteratively break down complex solutions into smaller, more manageable steps. Given a problem prompt $\boldsymbol{x}$, the algorithm outputs a decomposition of a solution, $\boldsymbol{y}=(\boldsymbol{y}_{1},\boldsymbol{y}_{2},\dots,\boldsymbol{y}_{K})$, such that the concatenation $\boldsymbol{y}_{1...K}$ forms a complete solution to $\boldsymbol{x}$.
The algorithm operates in two key stages:
1. Solution sampling. Starting from $\boldsymbol{x}$ and $\boldsymbol{y}_{0}=\emptyset$, the algorithm generates complete solutions $\boldsymbol{y}_{k}\sim\pi(\cdot|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(1-k)})$ using a policy $\pi$ like in single step generation. The best solution, $\boldsymbol{y}^{*}_{k}$, is selected based on the reward model $R$.
2. Recursive partitioning. The selected solution $\boldsymbol{y}^{*}_{k}$ is partitioned into two segments, $\boldsymbol{y}^{*}_{k}=\boldsymbol{y}^{*}_{a}\cdot\boldsymbol{y}^{*}_{b}$, based on a predefined partition fraction $\alpha$, where $|\boldsymbol{y}^{*}_{a}|\approx\alpha|\boldsymbol{y}^{*}_{k}|$. For each part, a priority metric $h$ is estimated: $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ and $\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, usually through rollouts of the step using $\pi$. The part with the lower priority is further partitioned.
  • If $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})\geq\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, additional samples are sampled for $\boldsymbol{y}^{*}_{b}$, with the process repeating on the new best solution, $\boldsymbol{y}^{\prime*}_{b}$.
  • Conversely, if $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})<\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, the first segment $\boldsymbol{y}^{*}_{a}$ is further partitioned. The first step corresponds to the $\alpha$ fraction of $\boldsymbol{y}^{\prime*}_{a}$, with the remaining part of the full solution forming the second step.
This recursive process is illustrated in Fig. 2. The pseudocode for DISC is provided in Alg. 1, with an annotated Python implementation in App. A.
The priority metric $h$ serves as the central heuristic for determining which solution steps to prioritize. It estimates the “difficulty” or “potential for improvement” of a step $\boldsymbol{y}_{k}$ given the context $\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)}$, computed via rollouts of the policy $\pi$. Specifically, $h(\boldsymbol{y}_{k}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ is estimated by sampling continuations and evaluating their outcomes.
In practice, estimating $h$ and generating new samples occur simultaneously, as both rely on rollout-based computations (Sec. 3.2). Unlike standard decomposition methods, DISC does not process steps in strict temporal order, resembling goal-directed planning (Parascandolo et al., 2020) and backtracking.
Intuition and Benefits. DISC partitions difficult steps into smaller, simpler sub-steps, allocating additional resources to refine them. This approach is particularly effective for inference scaling, where computing must be used judiciously. A binary decomposition strategy enables fast identification of difficult or high-potential steps.
Key Insight: Recursive partitioning Top-down, recursive partitioning means that we can both efficiently locate critical steps and also dynamically determine step sizes based on our budget.
Dynamic Compute Allocation. A key advantage of DISC is its ability to prioritize challenging or high-potential steps, improving solution quality while minimizing compute waste. By iteratively refining steps with low priority scores, DISC adaptively allocates more resources to difficult steps and less to simpler ones, optimizing inference efficiency.
Key Insight: Adaptive compute allocation DISC dynamically allocates inference compute to harder steps, optimizing solution quality and resource efficiency.
3.2 Priority Metric
We consider two intuitive priority metrics for step selection: Q-value priority and Z-score priority, which are visualized in Fig. 4.
Q-value based priority (DISC-Q). Given a partial solution $\boldsymbol{y}^{*}_{1...k}=\boldsymbol{y}^{*}_{1...(k-1)}\cdot\boldsymbol{y}^{*}_{a}\cdot\boldsymbol{y}^{*}_{b}$, we aim to prioritize either $\boldsymbol{y}^{*}_{a}$ or $\boldsymbol{y}^{*}_{b}$ for refinement. The core intuition behind Q-value prioritization is that steps with lower Q-values indicate areas needing refinement, directing compute toward the most challenging parts of the solution. More formally, we define the Q-priority metric under policy $\pi$ as:
$h_{Q}(\boldsymbol{y}^{*}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})=\mathbb{E}_{\boldsymbol{y}_{k}}\left[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})\right].$
Here, $\boldsymbol{y}_{k}$ represents alternative steps sampled from $\pi$, and $Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ denotes the Q-value of $\boldsymbol{y}_{k}$, conditioned on the partial solution $\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)}$. Equivalently, $h_{Q}(\boldsymbol{y}^{*}_{k})$ can be interpreted as the value function $V^{\pi}(\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$, where $V^{\pi}$ represents the expected reward achievable from the given partial solution. The expectation is taken over sampled candidates $\boldsymbol{y}_{k}\sim\pi(\cdot\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$, constrained by $|\boldsymbol{y}_{k}|=|\boldsymbol{y}^{*}_{k}|$. This metric helps identify difficult steps that the LLM is likely to get wrong, guiding partitioning toward the most critical refinements.
To estimate $h_{Q}$ for the second step $\boldsymbol{y}^{*}_{b}$, we sample $\boldsymbol{y}_{b}$ from $\pi$, generate rollouts $\boldsymbol{y}_{(k+1)+}$, compute rewards, and average the outcomes. For the previous step $\boldsymbol{y}^{*}_{a}$, we reuse rollouts from earlier partitioning, as the mean of these previously generated samples provides an unbiased estimate of the Q-priority metric:
$h_{Q}(\boldsymbol{y}^{*}_{a}\mid\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...(k-1)})=\mathbb{E}\left[R(\boldsymbol{y}^{*}_{1...(k-1)+})\right].$
By leveraging existing rollouts, DISC avoids redundant sampling, improving computational efficiency. Once the rollouts for $\boldsymbol{y}^{*}_{(k+1)}$ are available, the best completion $\boldsymbol{y}^{*}_{(k+1)+}$ is selected as the next step to partition. This dual use of rollouts optimizes both metric estimation and inference sampling.
Combining priority estimation and test-time search By integrating LLM policy sampling for both metric estimation and search, we can significantly enhance computational efficiency.
Figure 3: Priority metric estimation. We can estimate the priority metric $\hat{h}$ using Monte Carlo rollouts of the LLM policy $\pi$ for each step.
Z-score based priority (DISC-Z). To allocate more compute to steps with higher potential for improvement, we estimate the probability of sampling a better step given existing samples. Assuming the Q-values of sampled steps follow a normal distribution, we model this probability using the cumulative distribution function (CDF). Given a mean $\mu_{r}$ and standard deviation $\sigma_{r}$ of sampled Q-values, the probability of sampling better than the best-observed step $\boldsymbol{y}^{*}_{k}$ with Q-value $q^{*}$ is:
$1-\operatorname{CDF}\left(\frac{q^{*}-\mu_{r}}{\sigma_{r}}\right)=1-\operatorname{CDF}(z^{*}),$
where $z^{*}$ is the Z-score. Since the CDF is monotonic, we compare steps based on their Z-scores.
We formally define the Z-score priority metric as:
$h_{Z}(\boldsymbol{y}^{*}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})=\frac{q^{*}-\mathbb{E}[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})]}{\operatorname{Std}[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})]}.$
Since $q^{*}-\mu_{r}$ represents the advantage of step $\boldsymbol{y}^{*}_{k}$, the Z-score metric can be interpreted as a standard deviation-scaled advantage. Lower $h_{Z}$ values indicate steps with greater room for improvement, guiding decomposition toward those with higher variance in performance.
Figure 4: Different priority metrics. Sampled rewards of different steps can be visualized as a distribution.
3.3 DISC and Search Methods
DISC can also be used to enhance Monte Carlo Tree Search (MCTS) and other inference scaling and search methods. Recall that in our partition step, we greedily partition the best solution step $\boldsymbol{y}^{*}$. Instead of greedily partitioning the best step, we can partition the top $k$ best steps instead, and select which step to partition using the upper confidence tree (UCT) formula. We can also use beam search to prune out steps we do not want to partition further. We explain MCTS and beam search in detail in App. E and present results of combining DISC with search in Sec. 4.9.
Takeaway: Dynamic step sizes can improve search DISC can enhance search based inference scaling methods by determining what step size to search across.
3.4 A Motivating Example on DISC-Z
We use the Wiener process $W(t)$ as an example where there are intractably many actions and steps. Suppose we start at $t=0$ with $W(0)=0$. At each round $k$, the algorithm can choose one of the two options:
  1. 1.
    samples a trajectory and observe the final value $W(T)$ at time $t=T$, as the reward signal. Denote the whole trajectory as $w_{k}(\cdot)$.
  2. 2.
    chooses one trajectory from the previous rounds (denoted as $w_{s}(t)$ for round $s$), and time $t_{0}$; then sample a trajectory at $t=t_{0}$ with $W(t_{0})=w_{s}(t_{0})$. Denote the concatenated trajectory as $w_{k}(\cdot)$ with $w_{k}(t)=w_{s}(t)$ when $t\leq t_{0}$.
Note that we are only able to observe the final reward $W(t)$. At any intermediate time $t\in(0,T)$, the current value $W(t)$ is not observable. The goal is to design an algorithm that can reach the highest reward among the $K$ trajectories. Formally speaking, we aim to maximize the maximum:
$\max_{k\in K}w_{k}(T).$
One naive solution is to call option 1 for $K$ times and return the best-of-$K$ reward, each following:
$W(T)$ $\sim\mathcal{N}(0,T).$
Alternatively, suppose there is a promising path $w(\cdot)$ with a high final reward $w(T)=R$. It is natural to consider starting at some midpoint $\alpha T$ ($10<\alpha<$ 1) and perform more completions to obtain an even higher reward than $R$. The reward distribution sampled this way is
$W^{\prime}(T)\sim\mathcal{N}(w(\alpha T),(1-\alpha)T).$
The remaining question is which $\alpha$ we should choose. One option is to maximize the probability that the newly sampled reward is higher than $R$:
$\mathbb{P}(W^{\prime}(T)>R)=1-\Phi\bigg{(}\frac{R-w(\alpha T)}{\sqrt{(1-\alpha)T}}\bigg{)}.$
3.5 Example Decomposition
With sample budget 100, the decomposition of a MATH problem is as follows, where color indicates the value of the priority metric $h$ of each step (yellow low, purple high).
DISC example decomposition Let the length of the rectangle be $l$ and the width of the rectangle be $w$. Since the perimeter of the rectangle is 24 inches, we have that $242l+2w=$ 24, so $l+w=12$. We wish to maximize the area of the rectangle, which is $A=lw$. Let $l=12-w$ and plug into the area: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}A=(12-w)w\Rightarrow\qquad A=12w-w^{2}}$ Now, we differentiate $A$ with respect to $w$: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}A^{\prime}(w)=12-2w}$ We wish to maximize $A$, so we set $A^{\prime}(w)=0$, and solve for $w$: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}12-2w=0\Rightarrow\qquad w=6}$ Since $l=12-w$, we have that $l=12-6=6$. Therefore, the area of the rectangle is $A=lw=6\cdot 6=\boxed{36}$.
Once the LLM generates the first three steps, the rest is easy. Interestingly, ‘which’ is an important decision point which helps decide how the LLM will complete the solution.
Takeaway: Autoregressive models require autoregressive decomposition While words such as ‘which’, ‘therefore’, etc. may not seem like important steps to humans, they actually represent important steps for autoregressive LLMs which are trained on next token prediction.
4 Experimental Results
4.1 Benchmarks
We evaluate DISC on three benchmarks: APPS, MATH, and LiveCodeBench, to assess its impact on inference scaling for both coding and reasoning.
APPS (Hendrycks et al., 2021a) consists of 5000 competitive programming problems across three difficulty levels, with the competition-level subset being the hardest. We evaluate on a 200-problem subset due to computational constraints. MATH (Hendrycks et al., 2021b) comprises 12,500 math problems. Since the ground-truth verifier provides only binary rewards, we use a pretrained ORM (Xiong et al., 2024), trained via the method in (Wang et al., 2024b), with Llama-3.1-8B-Instruct as the base model. We test on a 500-problem subset (MATH500), identical to prior work (Wang et al., 2024b; Lightman et al., 2023). LiveCodeBench (Jain et al., 2024) is a continuously updated dataset from Leetcode, AtCoder, and CodeForces, ensuring LLMs have not been exposed to test problems. We evaluate on the 108 problems uploaded between 10/01/2024 and 12/01/2024 to prevent contamination.
Figure 5: Token-level comparison on APPS competition level. DISC consistently outperforms prior decomposition methods in efficiency.
Figure 6: Token-level comparison on MATH500. DISC outperforms other decomposition methods, demonstrating its effectiveness in mathematical reasoning.
Figure 7: Token-level comparison on LiveCodeBench. DISC achieves superior inference scaling over TokenSplit, LineSplit, and BoN.
4.2 Decomposition Comparison
We compare DISC against three prior decomposition methods from Sec. 2.2: TokenSplit (token-level decomposition), LineSplit (newline-based decomposition), and BoN (treating the entire solution as a single step).
Across all benchmarks, DISC achieves superior scaling and performance under both fixed token budgets (Fig. 7) and sample budgets (Fig. 12). We evaluate two key metrics: Pass@k, the proportion of problems solved within a sample budget $k$, and Pass@token, the proportion solved within a given token budget. Notably, DISC consistently outperforms static decomposition methods on APPS, MATH, and LiveCodeBench (Fig. 567), demonstrating its ability to allocate compute adaptively for improved inference efficiency. Extended results and analyses for each benchmark are provided in App. D.1D.4, and D.5.
4.3 Decomposition Analysis and Interpretation
Our results strongly indicate that decomposition—whether line-based, token-based, or DISC —improves sample quality. Figures 8 and 9 illustrate how the mean and variance of sampled rewards evolve with the step number, which represents the order in which a step is explored. Higher step numbers correspond to deeper search levels, where solutions are partitioned into finer-grained steps. As shown in Fig. 8, increasing step number correlates with higher-quality solutions, demonstrating that finer-grained decomposition improves sample quality. Additionally, Fig. 9 shows that reward variance decreases as step count increases, highlighting how decomposition enhances sampling precision.
Furthermore, DISC achieves better performance with fewer partitions under a fixed sampling budget (Fig. 10). We distinguish between actual partitions, the number of steps effectively explored, and planned partitions, the number of partitions intended by the method. Token and line split methods generate a large number of planned partitions (Fig. 10) but search over at most 15 steps due to budget constraints. In contrast, DISC dynamically adjusts the number of partitions based on available budget, efficiently identifying and focusing on critical steps.
Takeaway: Decomposition and sample quality Finer-grained decomposition improves sampled solution quality and reduces reward variance.
Figure 8: Average reward per step, averaged over APPS problems. From step 3 onward, higher step counts strongly correlate with increased average reward, demonstrating the effectiveness of decomposition. The dip between steps 1 and 3 likely occurs because simple problems are solved early, preventing further search.
Figure 9: Standard deviation of rewards per step, averaged over APPS problems. Decomposition reduces sampling variance, improving precision at deeper search depths.
Figure 10: Comparison of actual and planned partitions on APPS. DISC outperforms other methods with fewer partitions by efficiently identifying critical steps. Unlike token and line split methods, which plan many partitions but search only a subset, DISC dynamically adjusts partitioning based on budget.
4.4 Interaction Between Temperature and DISC
Figure 11: Effect of temperature on DISC performance on APPS with gpt-4o-mini. Unlike BoN and other inference scaling methods, DISC achieves higher performance at lower temperatures.
We perform ablation studies to analyze the impact of temperature on DISC. Typically, inference scaling methods achieve optimal performance at temperatures around 0.6–0.8, as increased temperature promotes sample diversity (Wang et al., 2024a). Surprisingly, however, DISC performs better at lower temperatures, as shown in Fig. 11. This trend is in stark contrast to BoN (Fig. 22), where higher temperatures are generally beneficial.
We believe this phenomenon arises because DISC depends on accurately estimating the priority metric $h$ at each step. Lower temperatures reduce sample variance, leading to more reliable estimates of $h$, which in turn improves step selection. This is further supported by Fig. 20, which shows that lower temperatures yield lower standard deviations per step, indicating increased sampling consistency. Additional details and analyses can be found in App. C.1.
4.5 Self-Generated Validation Tests
We also evaluate DISC in a more practical setting where a ground-truth reward model is unavailable for code generation  (Chen et al., 2022, 2023b; Zhou et al., 2024). Instead of relying on predefined test cases, we prompt the LLM to generate validation test cases based on the problem prompt. In real-world applications, manually curated ground-truth test cases are often costly to obtain, making self-generated validation a more scalable approach. The results, shown in Fig. 12, indicate that DISC continues to scale better than other methods in this setting. Additional results and details are provided in App. D.3.
Figure 12: Comparison of Pass@k performance on APPS with ground truth tests (right) and self generated validation tests (left) using gpt-4o-mini. DISC scales more effectively in both.
4.6 Ablation on Base LLM Model
We evaluate DISC across different LLMs, including open-source models. As shown in Fig. 13 and Fig. 27, DISC significantly enhances performance even for weaker models. Specifically, it improves Llama’s pass rate from 1% to 5.5%, a 550% relative increase, and Mistral’s from 0% to 3.5%, demonstrating substantial gains even from a nonzero baseline. Additional details and analyses are in App. C.3.
Figure 13: Pass@k scaling curve for open-source models with DISC on APPS. DISC substantially improves performance across different LLMs, including Llama and Mistral.
4.7 Interaction Between Priority Metric and DISC
We conduct an ablation study to examine how the choice of priority metric affects DISC performance. In addition to the Q-based and Z-based priority metrics (DISC-Q and DISC-Z) introduced in Sec. 3.2, we evaluate three baselines: DISC-R (random step selection), DISC-negQ, and DISC-negZ (which prioritize the opposite steps of DISC-Q and DISC-Z, respectively). As shown in Fig. 14, the selection of a priority metric significantly impacts performance. Both DISC-Q and DISC-Z significantly outperform random selection and their inverse counterparts, demonstrating the effectiveness of their priority heuristics. Additional details and analysis are in App. C.2.
Figure 14: Effect of priority metrics on DISC performance on APPS with gpt-4o-mini. Both Q-based and Z-based priority metrics outperform random selection and their inverses, highlighting their effectiveness.
4.8 Ablation on Partition Fraction $\alpha$
We conduct an ablation study to analyze the effect of the partition fraction $\alpha$ on DISC performance. As shown in Fig.  15 and  31, the optimal range appears to be $0.15\leq\alpha\leq 0.25$. Lower partition fractions ($\alpha<0.5$) tend to perform better due to the asymmetric cost of sampling from different halves of the partition. Sampling from the first half requires generating more tokens, while the second half requires fewer, making it crucial to partition the first half more conservatively. Additional analysis are in App. C.4.
Figure 15: Pass@k comparison of different DISC partition fractions $\alpha$ on APPS competition level. The range $0.15\leq\alpha\leq 0.25$ appears optimal.
4.9 Search and DISC
We demonstrate that search methods such as MCTS and beam search can be combined with DISC. As shown in Fig. 47 in the Appendix, greedy search explores deeper partitions given the same search budget due to its greedy nature, while MCTS and beam search reach similar, shallower depths. However, MCTS allocates the search budget more effectively than beam search, leading to higher performance, as seen in Fig. 16. Additional details and analysis are in App. E.
Figure 16: Pass@k comparison of search methods combined with DISC on APPS with gpt-4o-mini. MCTS scales best, followed by greedy search, then beam search (with beam size 2).
5 Related Work
Inference scaling. Inference scaling has emerged as a dominant paradigm, driven by the introduction of o1- and r1-like chain-of-thought reasoning models (Snell et al., 2024; Brown et al., 2024; Manvi et al., 2024; Leea et al., 2025). Several works examine the trade-off between inference compute and training compute (Guan et al., 2025; Chen et al., 2024b). LLM inference often relies on decomposing complex problems into intermediate reasoning steps, as seen in chain-of-thought (CoT) prompting (Wei et al., 2022; Sprague et al., 2024; Wang & Zhou, 2024) and its variants (Kojima et al., 2022; Zhou et al., 2023; Wang et al., 2023b; Li et al., 2023). We extend inference scaling by introducing a new approach for adaptive compute allocation (Manvi et al., 2024).
LLM reasoning and code generation. LLM reasoning and code generation are central tasks for inference scaling. Evolutionary inference scaling methods have been explored in program generation (Liventsev et al., 2023; Chen et al., 2023a; Romera-Paredes et al., 2024; Lehman et al., 2023; Hemberg et al., 2024). Domain-specific decomposition strategies have been applied in code generation, such as function-based decomposition (Chen et al., 2024a; Zenkner et al., 2024; Levin et al., 2025). More broadly, decomposition often involves prompting LLMs to generate subtask completions (Hernández-Gutiérrez et al., 2024; Khot et al., 2022; Dua et al., 2022), which differs from methods that refine a single LLM generation.
Reinforcement learning and Monte Carlo methods. Unlike standard RL, our setting resembles a search problem where the goal is to identify the single highest-reward path. Cazenave (2009) demonstrated that nested Monte Carlo search can accelerate optimal pathfinding. Under the bandit setting, this can be formulated as identifying the arm with the highest maximum reward rather than the highest mean reward (Cicirello & Smith, 2005; Carpentier & Valko, 2014).
6 Conclusion
We introduce DISC, a dynamic decomposition framework that adaptively partitions solution steps based on difficulty, improving inference scaling by directing compute toward critical steps while balancing exploration and resource allocation. DISC seamlessly integrates with search-based methods such as MCTS and beam search, further enhancing performance. It also identifies challenging steps for LLMs, aiding curriculum learning, fine-tuning, and dataset augmentation. By dynamically adjusting partitioning based on available compute, DISC enables more adaptive and efficient reasoning in large language models, with broad implications for both training and inference optimization.
7 Impact Statements
This work on dynamic decomposition aims to advance the field of LLM by improving the efficiency and accuracy of large-scale inference methods, particularly in domains like code generation, mathematical problem-solving, and automated reasoning. By dynamically allocating computational resources to challenging reasoning steps, this approach has the potential to enhance AI-powered tools in software development, STEM education, and scientific research. While the core contributions of this paper focus on technical advancements, we acknowledge the broader societal implications of scaling inference techniques, including risks of over-reliance on AI in decision-making and the potential propagation of errors or biases. To mitigate these concerns, future work should emphasize robust validation, fairness, and human-in-the-loop deployment strategies, particularly in high-stakes applications. Nonetheless, no immediate or specific ethical concerns are identified in the context of this research.
Appendix A Code implementation of DISC
Python implementation of DISC def dynamic_decomposition(problem, model, reward_model, split_str, complete_solution, fraction, solution_budget, split_metric, stop_threshold=-float("inf"), stop_sum_score=1.0, stop_if_solved=False, ): """ ␣␣␣␣Decomposesthesolutionusingadynamicbinarysearchapproach ␣␣␣␣Args: ␣␣␣␣␣␣␣␣problem(Problem):Theproblemtosolve ␣␣␣␣␣␣␣␣model(Model):Themodeltouseforgeneration ␣␣␣␣␣␣␣␣reward_model(function):Therewardmodeltouseforscoring ␣␣␣␣␣␣␣␣split_str(function):Thefunctiontouseforsplittingastring ␣␣␣␣␣␣␣␣complete_solution(function):Thefunctiontouseforcompletingasolution ␣␣␣␣␣␣␣␣fraction(float):Thefractiontosplitthestring ␣␣␣␣␣␣␣␣solution_budget(int):Themaximumnumberofsolutionstogenerate ␣␣␣␣␣␣␣␣split_metric(function):Themetrictouseforsplitting ␣␣␣␣␣␣␣␣stop_threshold(float):Thethresholdtostopsplitting ␣␣␣␣␣␣␣␣stop_sum_score(float):Thesumscoretostopgeneratingcompletions ␣␣␣␣␣␣␣␣stop_if_solved(bool):Whethertostopiftheproblemissolved ␣␣␣␣""" # Initialize results and decomposition steps decomp_return = { "generated_solutions": [], "decomposition": [] } while len(decomp_return["generated_solutions"]) < solution_budget: # Combine all previous steps into an intermediate solution intermediate_solution = "".join([step["step_str"] for step in decomp_return["decomposition"]]) new_scores = [] best_solution = None best_completion = None best_score = -float("inf") sum_score = 0.0 # 1) Generate completions until we generate enough samples to estimate the split metric while sum_score < stop_sum_score: proposed_completion = complete_solution(problem, intermediate_solution, model) proposed_solution = intermediate_solution + proposed_completion decomp_return["generated_solutions"].append(proposed_solution) # Update scores proposed_score = reward_model(proposed_solution) new_scores.append(proposed_score) sum_score += proposed_score # Track the best solution if proposed_score > best_score: best_solution = proposed_solution best_score = proposed_score best_completion = proposed_completion # Stop early if problem is solved if stop_if_solved and proposed_score >= 1.0: decomp_return["decomposition"].append({"step_str": proposed_completion}) return decomp_return new_metric = split_metric(new_scores) last_metric = decomp_return["decomposition"][-1]["metric"] if decomp_return["decomposition"] else None # Determine the split target. We always split the step with the highest metric is_split_new_step = last_metric is None or new_metric >= last_metric split_target = decomp_return["decomposition"][-1]["step_str"] if not is_split_new_step else best_completion # 3) Attempt to split the target split_result = split_str(split_target, fraction) if not split_result: # If we can’t split the target, we’re done decomp_return["decomposition"].append({"step_str": best_completion, "metric": new_metric}) return decomp_return # Update decomposition based on split part1, part2 = split_result if is_split_new_step: decomp_return["decomposition"].append({"step_str": part1, "metric": new_metric}) # Stopping condition based on threshold if new_metric < stop_threshold: decomp_return["decomposition"].append({"step_str": part2}) return decomp_return else: decomp_return["decomposition"][-1] = {"step_str": part1, "metric": last_metric} return decomp_return
Appendix B Pseudocode for DISC
Algorithm 1 Dynamic Decomposition
1:  Input: Problem instance ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}x}$, reward model ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r}$, partition function ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}f}$, LLM policy model ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\pi}$, partition fraction ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\alpha}$, solution budget ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$, priority metric ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}h}$, metric stopping precision ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$, sampling stopping threshold ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\sigma}$, is inference mode binference
2:  Output: Final decomposition ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$
[…]
4:  # Decompose the solution recursively until we reach the desired precision ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$ or run out of budget ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$
5:  while $|{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{generated\_solutions}|<{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$ do
6:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}}\leftarrow\text{Concatenate}([\text{step.step\_str}\ \forall\ \text{step}\in{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}])$
[…]
8:     ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}}\leftarrow\varnothing$vartriangleright Record rewards of completions
9:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{final}}}\leftarrow\text{None}$, ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\leftarrow\text{None}$, ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}\leftarrow-\infty$vartriangleright Track the best completion
10:     # Step 1: Generate completions until we have enough samples to estimate the splitting metric. Here we use a geometric sampling distribution
11:     while $\operatorname{sum}({\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}})<{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\sigma}$ do
12:        ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\pi}(\cdot|{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}x},{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}})$
13:        ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}}\oplus{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}$
[…]
15:        ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}\leftarrow{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r}({\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}})$
16:        Append ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}$ to ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}}$
17:        if ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}>{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}$ then
18:           ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{final}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}}$, ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}$
19:           ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}\leftarrow{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}$
20:        end if
21:        if $\textbf{b}_{\text{inference}}\ \textbf{and}\ {\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}=1.0$ then
[…]
23:           Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if problem is solved
24:        end if
25:     end while
26:     # Step 2: Compute splitting metric
27:     ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\leftarrow{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}h}({\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}})$
28:     ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}[-1].{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z}\ \textbf{if}\ {\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}\neq\varnothing\ \textbf{else}\ -\infty$
29:     # Step 3: Split the step with the higher metric
30:     bsplit new step $\leftarrow$ ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\geq{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}$
[…]
32:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}},{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}f}({\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{target step}}},{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\alpha})$
33:     if ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}}=\text{None}\ \textbf{or}\ {\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}=\text{None}$ then
[…]
35:        Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if we cannot do a finer split
36:     end if
37:     if $\textbf{b}_{\text{split new step}}$ then
[…]
39:        if ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}<{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$ then
[…]
41:           Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if all metrics are smaller than precision
42:        end if
43:     else
44:        ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}[-1]\leftarrow\{\text{step\_str}:{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}},\text{metric}:{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}\}$vartriangleright Split last step
45:     end if
46:  end while
47:  Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$
Appendix C Ablation studies
C.1 Ablation on Temperature
We conduct an ablation study to analyze the effects of temperature on DISC and BoN. Temperature controls the randomness of token sampling in autoregressive models, influencing both exploration and consistency. Higher temperatures encourage more diverse outputs, whereas lower temperatures yield more deterministic generations. To examine its impact, we evaluate DISC and BoN on a 100-problem subset of APPS (the first 100 problems) using gpt-4o-mini.
Figure 17 presents the Pass@token scaling curve for DISC across different temperatures. The results indicate that lower temperatures lead to improved performance, as DISC benefits from more deterministic step selection. Unlike BoN, which relies on broad solution sampling, DISC dynamically refines steps, making stable token probabilities advantageous.
Figure 18 illustrates the frequency of actual partitions made by DISC at different temperatures. As temperature increases, the number of partitions fluctuates more, suggesting that high temperature introduces instability in step selection. Lower temperatures provide more structured decomposition, reducing unnecessary subdivisions.
In Figure 19, we visualize the mean reward per step. The trend shows a linear increase in reward as step number grows, demonstrating that deeper decomposition results in progressively better solutions. This reinforces that DISC effectively allocates computation towards refining difficult steps.
The mean standard deviation per step is shown in Figure 20. Lower temperatures yield lower standard deviations, confirming that DISC benefits from reduced variability in sample quality. This consistency allows for more reliable prioritization of difficult steps, enhancing overall inference efficiency.
For comparison, Figure 22 and Figure 21 display Pass@token and Pass@k scaling curves for BoN across different temperatures. Unlike DISC, BoN achieves peak performance at a temperature around 0.6-0.8, balancing diversity and consistency. Higher temperatures increase exploration but degrade precision, while lower temperatures hinder sample diversity, reducing the probability of obtaining high-quality completions.
These findings highlight the fundamental difference between DISC and BoN: DISC benefits from lower variance and stable decomposition, while BoN relies on broader exploration facilitated by moderate temperature settings. As a result, optimal temperature settings differ significantly between these methods, with DISC favoring deterministic sampling and BoN requiring a balance between diversity and coherence.
Figure 17: Pass@token scaling curve for different temperatures on APPS using gpt-4o-mini. The lower the temperature, the stronger the DISC performance.
Figure 18: Partition frequency of DISC with different temperatures on APPS using gpt-4o-mini
Figure 19: Mean reward per step of DISC with different temperatures on APPS using gpt-4o-mini. The mean reward scales linearly with step number.
Figure 20: Mean standard deviation per step of DISC with different temperatures on APPS using gpt-4o-mini. Lower temperature means lower average standard deviation.
Figure 21: Pass@k scaling curve for different temperatures on APPS using gpt-4o-mini for BoN. A temperature around 0.6-0.8 leads to the best performance and balance between diversity and consistency.
Figure 22: Pass@token scaling curve for different temperatures on APPS using gpt-4o-mini for BoN. A temperature around 0.6-0.8 leads to the best performance and balance between diversity and consistency.
C.2 Ablation on Priority Metric $h$
We analyze the effect of different priority metrics on DISC performance. We evaluate DISC using the first 200 competition-level APPS problems with gpt-4o-mini, setting the temperature to 0.8 for all experiments. The priority metric determines which steps are refined during recursive decomposition, impacting both efficiency and final solution quality.
Figure 23 presents a token-level comparison of different priority metrics. Both DISC-Q and DISC-Z significantly outperform random selection and their inverse counterparts, demonstrating the importance of prioritizing high-value steps.
Figure 24 illustrates the partition frequency under different priority metrics. We observe that effective metrics such as DISC-Q and DISC-Z lead to fewer, more meaningful partitions, whereas suboptimal strategies result in excessive, redundant partitioning.
The relationship between mean reward and step number is shown in Figure 25. All tested metrics exhibit a strong correlation between increasing step depth and mean reward, indicating that decomposition progressively refines solutions. However, DISC-Q and DISC-Z achieve higher reward gains at earlier stages, suggesting that they prioritize the most impactful refinements.
Finally, Figure 26 reports the standard deviation of rewards per step. Lower standard deviation suggests more stable solution quality, a property that DISC-Q and DISC-Z maintain better than random selection methods. This highlights their effectiveness in identifying and refining challenging steps efficiently.
Overall, these results confirm that choosing an appropriate priority metric is crucial for DISC. While DISC-Q and DISC-Z consistently enhance inference efficiency and quality, random or inverse strategies lead to poorer performance due to misallocation of compute resources.
Figure 23: Token level comparison of different priority metrics on DISC in the APPS setting with gpt-4o-mini. Both Q and Z based priority metrics perform well.
Figure 24: Partition frequency of DISC with different priority metrics on APPS using gpt-4o-mini
Figure 25: Mean reward per step of DISC with different priority metrics on APPS using gpt-4o-mini. All metrics display strong correlation between step depth and the mean reward.
Figure 26: Mean standard deviation per step of DISC with different priority metrics on APPS using gpt-4o-mini. All metrics display correlation between step depth and the standard deviation.
C.3 Model Ablation
We investigate how different LLMs perform when used with DISC on 200 competition-level APPS problems, given a sample budget of 30. The groundtruth reward model was used to evaluate correctness, and all models were set to a temperature of 0.8. Due to the challenging nature of the benchmark, open-source models struggled to achieve strong performance independently. However, when paired with DISC, their performance significantly improved.
Figure 27 presents the Pass@token scaling curve for open-source models using DISC. The results demonstrate that DISC substantially enhances the capabilities of these models, closing the gap between them and proprietary alternatives.
Figure 28 visualizes the partition frequency of DISC with different open-source models. Compared to their standalone performance, the use of DISC led to more structured and effective decomposition, highlighting its adaptability to different architectures.
The mean reward per step is shown in Figure 29. Similar to prior findings, we observe that deeper decomposition leads to increasingly higher rewards. Notably, even lower-capacity models benefit from DISC ’s ability to iteratively refine their solutions.
Finally, Figure 30 presents the mean standard deviation per step. With DISC, the variance in performance is significantly reduced, resulting in more stable and reliable inference.
Overall, these findings emphasize that DISC is a robust framework capable of enhancing inference performance across diverse LLMs, particularly those with limited standalone capabilities.
Figure 27: Pass@token scaling curve for open source models with DISC on APPS. DISC also demonstrates strong performance gains with open source models.
Figure 28: Partition frequency of DISC with open source models on APPS
Figure 29: Mean reward per step of DISC with open source models on APPS
Figure 30: Mean standard deviation per step of DISC with open source models on APPS
C.4 Ablation on partition fraction $\alpha$
We include some more analysis on the partition fraction here.
Figure 31: Token level comparison of different DISC splitting fraction $\alpha$ on APPS competition level. $0.15\leq\alpha\leq 0.25$ seems to be optimal.
Appendix D Experimental results extended
D.1 APPS
Figure 32: Sampling frequency of each step averaged over the problems on APPS with gpt-4o-mini. DISC seems to have a preferences for spending more compute on earlier found steps.
D.2 Additional examples
Below is another computed decomposition for the same problem as the one shown in the main text:
Let $x$ be the length and $y$ be the width of the rectangle. Since the perimeter is 24 inches, we have $122x+2y=24\Rightarrow x+y=$ 12 . Therefore, we must maximize the area $xy$ subject to this constraint.
We use the method of Lagrange multipliers. The Lagrangian is
${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}\mathcal{L}(x,y,\lambda)=xy-\lambda(x+y-12).}$
We differentiate with respect to $x,y,$ and $\lambda$ to obtain
${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}\frac{\partial\mathcal{L}}{\partial x}=y-\lambda=0,\qquad\frac{\partial\mathcal{L}}{\partial y}=x-\lambda=0,\qquad\frac{\partial\mathcal{L}}{\partial\lambda}=x+y-12=0.}$
We find that $x=y=\lambda$, so $x=y=\frac{12}{2}=6.$ Therefore, the maximum area of the rectangle is $6\cdot 6=\boxed{36}.$
D.3 APPS with self-generated validation tests
We examine DISC performance on APPS when using self-generated validation tests. All methods utilized the same set of self-generated validation tests to ensure fair comparisons. Each problem received 5-10 validation tests, with the exact number determined dynamically by the LLM. We evaluated a subset of 100 APPS problems, generating samples until the sample budget was exhausted or a correct solution was found.
Figure 33 illustrates the Pass@token scaling curve, showing that DISC maintains strong scaling performance in this setting, though at a slightly lower rate compared to ground-truth verification.
Figure 34 and Figure 35 compare actual and planned partition frequencies, respectively. The results indicate that DISC continues to make structured decompositions even with self-generated validation, preserving its efficiency.
The mean reward per step, shown in Figure 36, follows a similar trend as in previous experiments, reinforcing that DISC effectively allocates compute resources for iterative refinement.
Lastly, Figure 37 demonstrates that DISC maintains lower standard deviations in performance, indicating stable quality improvements across steps.
Figure 33: Token level comparison of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests. DISC still scales better than other methods in this setting, albeit at a lower rate.
Figure 34: Actual partition frequency of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
Figure 35: Planned partition frequency of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
Figure 36: Mean reward per step of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
Figure 37: Mean standard deviation different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
D.4 MATH500
Completions for MATH500 include both the reasoning steps and the final answer. Since MATH500 contains more problems than APPS200 and MATH problems tend to be relatively easier, solution quality saturates quickly. Therefore, we use a lower sample budget of 10 for these experiments.
Figure 38 presents the Pass@k performance for different decomposition methods on MATH500. We observe that all decomposition-based approaches achieve similar Pass@k performance, consistently outperforming BoN. This indicates that the structured nature of MATH problems allows multiple decomposition strategies to be effective.
Despite similar Pass@k results, the true advantage of DISC lies in its token efficiency, as shown in Figure 6. DISC significantly reduces the number of tokens required to reach correct solutions compared to alternative methods, demonstrating its ability to allocate computational effort efficiently in mathematical reasoning tasks.
Additionally, we analyze the partitioning behavior of DISC on MATH500. Figure 39 illustrates the actual partition frequency for different decomposition methods. The planned partitioning behavior, shown in Figure 40, further highlights how DISC effectively balances exploration and refinement.
Finally, we present the mean standard deviation per step in Figure 41. Lower variance suggests that DISC produces more stable and reliable decompositions over multiple runs, reinforcing its robustness in both mathematical and program synthesis domains.
Figure 38: Pass@k performance comparison for different decomposition methods on MATH500. DISC consistently outperforms BoN across different sampling budgets.
Figure 39: Observed partition frequency of different decomposition methods on MATH500. DISC effectively segments problems into meaningful subcomponents.
Figure 40: Planned partitioning strategy of different decomposition methods on MATH500. DISC ’s structured approach leads to more efficient problem breakdowns.
Figure 41: Mean standard deviation per step for different decomposition methods on MATH500. Lower variance in DISC suggests more stable and reliable problem-solving steps.
D.5 LiveCodeBench
We evaluate DISC on LiveCodeBench, a benchmark designed for code generation tasks with a focus on real-world software development challenges. LiveCodeBench presents a unique set of problems requiring both reasoning and structured decomposition, making it a suitable testbed for evaluating DISC’s ability to refine and improve intermediate steps.
Figure 42 shows the Pass@k comparison of different decomposition methods on LiveCodeBench. DISC consistently scales better than other decomposition methods, highlighting its ability to refine intermediate steps more effectively in complex coding scenarios.
Figure 43 illustrates the observed partition frequency of different decomposition methods. The structured approach of DISC results in well-balanced decomposition across steps, reducing unnecessary partitioning while maintaining sufficient granularity for improved solution refinement.
Figure 44 displays the planned partition frequency across methods. DISC dynamically determines the most effective partitions based on the evolving problem state, leading to more targeted and efficient decompositions.
Finally, Figure 45 presents the mean standard deviation per step across decomposition methods. Lower variance in DISC suggests that it produces more stable and reliable decompositions, reinforcing its robustness for solving LiveCodeBench problems.
Figure 42: Pass@k performance comparison for different decomposition methods on LiveCodeBench. DISC consistently outperforms other methods in structured problem refinement.
Figure 43: Observed partition frequency of different decomposition methods on LiveCodeBench. DISC effectively balances problem segmentation while avoiding excessive partitioning.
Figure 44: Planned partitioning strategy of different decomposition methods on LiveCodeBench. DISC dynamically adapts its partitioning to optimize search efficiency.
Figure 45: Mean standard deviation per step for different decomposition methods on LiveCodeBench. Lower variance in DISC suggests more stable and reliable problem-solving steps.
Appendix E Search and scaling
E.1 Monte Carlo Tree Search (MCTS)
Monte Carlo Tree Search (MCTS) is a widely used algorithm for sequential decision-making in large search spaces, particularly in applications such as game playing, planning, and inference scaling. The algorithm builds a search tree incrementally by simulating different sequences of actions and updating estimates of state quality. A key advantage of MCTS is its ability to balance exploration (discovering new states) and exploitation (refining promising ones) using a data-driven search process. The MCTS pipeline consists of four fundamental steps: selection, expansion, simulation, and backpropagation.
E.1.1 Selection
Starting from the root node representing the current state $\boldsymbol{s}$, MCTS iteratively traverses the search tree by selecting child nodes based on a selection policy. The most commonly used selection criterion is the Upper Confidence Bound for Trees (UCT), which balances exploration and exploitation:
$UCT(\boldsymbol{s},\boldsymbol{d})=\hat{Q}(\boldsymbol{s},\boldsymbol{d})+c\sqrt{\frac{\ln\left(\sum_{\boldsymbol{b}}n(\boldsymbol{s},\boldsymbol{b})\right)}{n(\boldsymbol{s},\boldsymbol{d})}},$ (1)
where $\hat{Q}(\boldsymbol{s},\boldsymbol{d})$ represents the estimated value of selecting action $\boldsymbol{d}$ from state $\boldsymbol{s}$, $n(\boldsymbol{s},\boldsymbol{d})$ is the visit count for this action, and $c$ is a hyperparameter controlling the trade-off between exploring new actions and favoring those with high past rewards.
E.1.2 Expansion
Once a leaf node (a previously unexplored state) is reached, the algorithm expands the tree by adding one or more new nodes. These new nodes represent potential future states $\boldsymbol{s}^{\prime}$ generated by sampling an action $\boldsymbol{d}$ from a predefined policy. This step broadens the search space and allows MCTS to evaluate new possibilities.
E.1.3 Simulation
Following expansion, the algorithm conducts a simulation (or rollout) from the newly added state. This step involves generating a sequence of actions according to a predefined policy until reaching a terminal state or an evaluation horizon. The outcome of the simulation, denoted as $v(\boldsymbol{s}^{\prime})$, provides an estimate of the quality of the new state. Depending on the application, this could represent a game result, an optimization score, or an inference accuracy metric.
E.1.4 Backpropagation
The final step involves propagating the results of the simulation back up the search tree to refine the estimated values of prior states and actions. Each node along the trajectory $\tau=[\boldsymbol{s}_{0},\boldsymbol{d}_{1},\boldsymbol{s}_{2},\dots,\boldsymbol{s}_{-1}]$ is updated iteratively:
$\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t+1)}\leftarrow(1-\alpha_{n})\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t)}+\alpha_{n}\max\{\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t)},\hat{Q}(\boldsymbol{s}_{i+1},\boldsymbol{d}_{i+2})^{(t+1)}\},$ (2)
where $\alpha_{n}$ is a learning rate that depends on the visit count, and the maximum function ensures that the best-performing trajectories are emphasized.
MCTS has been widely adopted in inference scaling techniques due to its ability to efficiently allocate computational resources, focusing more on high-reward states while avoiding unnecessary exploration of unpromising regions. In later sections, we explore how MCTS can be combined with dynamic decomposition to further optimize inference scaling.
E.1.5 Combining Dynamic Decomposition with MCTS
MCTS can be enhanced by integrating dynamic decomposition, where each node in the search tree represents a decomposition of the problem into steps. Instead of treating states as atomic decisions, we recursively decompose reasoning steps, dynamically adjusting granularity based on difficulty.
In this framework:
  • Each node in the MCTS tree represents a partial decomposition of the problem, with child nodes corresponding to alternative step partitions.
  • Branching occurs by generating candidate next steps using dynamic decomposition, allowing finer steps for complex regions while maintaining efficiency for simpler ones.
  • The selection step prioritizes nodes that represent more promising decompositions, dynamically refining challenging areas through recursive subdivision.
  • The backpropagation step ensures that decompositions leading to high-quality solutions are reinforced, helping the search tree converge toward optimal inference paths.
By integrating dynamic decomposition with MCTS, we efficiently allocate compute to the most critical reasoning steps, improving inference quality while maintaining computational efficiency.
E.2 Beam Search
Beam search is a heuristic search algorithm commonly used in inference tasks where computational efficiency is a priority. Unlike exhaustive search methods, beam search maintains only the top $k$ best candidates at each step, making it an effective strategy for structured prediction problems and sequential decision-making.
At each iteration:
  • The algorithm selects the $k$ most promising partitions from the previous step based on an evaluation metric.
  • Each selected partition is expanded by generating possible next-step samples.
  • The newly generated partitions are ranked, and only the top $k$ candidates are retained for the next iteration.
  • This process continues until a stopping criterion is met, such as reaching a predefined depth or finding a sufficiently high-quality solution.
Beam search provides a computationally efficient way to explore structured solution spaces while maintaining high-quality search trajectories. By integrating beam search with dynamic decomposition, we ensure that inference computation is allocated efficiently, focusing on the most promising reasoning paths at each step.
E.3 Additional Results and Analysis
Experiments comparing different search methods were conducted on a 100-problem subset of the APPS dataset (first 100 problems) using GPT-4o-mini. All methods used a temperature of 0.2, with $\alpha=0.15$, Q priority metric, and $\sigma=1.0$.
Token-level comparison: As shown in Figure 46, MCTS scales best among the tested methods, demonstrating superior efficiency in identifying promising partitions. Greedy search follows closely, while beam search exhibits the slowest scaling.
Partition frequency analysis: Figure 47 reveals that greedy search explores to greater depths within the same sampling budget. This suggests that greedy search prioritizes deep refinements, whereas MCTS and beam search balance depth with breadth.
Step variance analysis: Figure 48 illustrates that all search methods display decreasing standard deviation with increasing search depth. This trend indicates that deeper searches converge towards stable, high-quality partitions, reinforcing the benefits of dynamic decomposition.
These results highlight the trade-offs between search methods: MCTS offers robust exploration-exploitation balance, greedy search favors depth-first refinement, and beam search provides a structured yet computationally constrained approach. The integration of dynamic decomposition further enhances these search strategies by adaptively allocating computational resources to critical reasoning steps.
Figure 46: Token level comparison of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. MCTS scales best, followed by greedy search, followed by beam search.
Figure 47: Actual partition frequency of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. Greedy is able to search to higher depths given the same sampling budget.
Figure 48: Mean standard deviation of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. All search methods display decreasing standard deviation with search depth.

### v3

@@
 Abstract
-Inference scaling methods often rely on decomposing problems into steps (or groups of tokens), followed by sampling and selecting the best next steps.
+Inference scaling methods for LLMs often rely on decomposing problems into steps (or groups of tokens), followed by sampling and selecting the best next steps.
 However, these steps and their sizes are often predetermined or manually designed based on domain knowledge.
-We propose dynamic decomposition, a method that adaptively and automatically fractions solution and reasoning traces into manageable steps during inference.
+We propose dynamic decomposition, a method that adaptively and automatically partitions solution and reasoning traces into manageable steps during inference.
 By more effectively allocating compute – particularly through subdividing challenging steps and prioritizing their sampling – dynamic decomposition significantly improves inference efficiency.
-Experiments on benchmarks such as APPS, MATH, and LiveCodeBench demonstrate that dynamic decomposition outperforms static approaches, including token-level, sentence-level, and single-step decompositions.
+Experiments on benchmarks such as APPS, MATH, and LiveCodeBench demonstrate that dynamic decomposition outperforms static approaches, including token-level, sentence-level, and single-step decompositions, reducing the pass@10 error rate by 5.0%, 6.7%, and 10.5% respectively.
 These findings highlight the potential of dynamic decomposition to improve a wide range of inference scaling techniques.
-Machine Learning, LLMs, Inference scaling, test-time compute, planning, reasoning, decomposition
 1 Introduction
 Scaling inference efficiency remains a fundamental challenge for large language models (LLMs).
-Many existing approaches improve inference by decomposing problems into smaller steps and systematically exploring different solutions (Feng et al., 2023; Zeng et al., 2024; Wu et al., 2024; Nori et al., 2024; Snell et al., 2024; Brown et al., 2024; Gandhi et al., 2024; Lee et al., 2025; Light et al., 2024a; Wang et al., 2025).
-Some decomposition methods rely on domain-specific heuristics and hand-crafted rules (Yao et al., 2024; Zelikman et al., 2023; Zhou et al., 2022).
+Many existing approaches improve inference by decomposing problems into smaller steps and systematically exploring different solutions (feng2023alphazero, ; zeng2024scaling, ; wu2024inferencescalinglawsempirical, ; nori2024medprompto1explorationruntime, ; snell2024scaling, ; brown2024large, ; gandhi2024stream, ; lee2025evolvingdeeperllmthinking, ; light2024strategist, ; anonymous2025planning, ).
+Some decomposition methods often rely on domain-specific heuristics and hand-crafted rules (yao2024tree, ; zelikman2023parsel, ; zhou2022least, ).
 However, manually partitioning problems or designing task-specific heuristics is costly and lacks generalization.
 Moreover, identifying critical steps for an LLM can be non-trivial for humans.
-As shown in Sec.
-3.5, LLMs may assign importance to seemingly trivial words (e.g., therefore or which), which, while counterintuitive to humans, play a crucial role in autoregressive generation (Lin et al., 2025).
-Other approaches employ fixed, uniform step sizes, such as token- or sentence-level decomposition (Feng et al., 2023; Guo et al., 2025).
+LLMs may assign importance to seemingly trivial words (e.g., therefore or which), which, while counterintuitive to humans, play a crucial role in autoregressive generation (lin2025criticaltokens, ).
+Other approaches employ fixed, uniform step sizes, such as token- or sentence-level decomposition (feng2023alphazero, ; guo2025deepseek, ).
 All these methods rely on static decomposition strategies, where step sizes are predefined or determined via heuristics.
-Such rigidity wastes compute on steps that are easy for the LLM (but potentially difficult for humans) while undersampling more challenging steps.
+Such rigidity risks overusing compute on steps that are easy for the LLM (but potentially difficult for humans) while undersampling more challenging steps.
 To overcome these limitations, we propose DISC (Dynamic decomposition Improves Scaling Compute), a recursive inference algorithm that dynamically partitions solution steps based on difficulty.
 Unlike prior methods, DISC adapts decomposition granularity during inference based on both the available budget and problem complexity, ensuring finer granularity for more difficult steps.
-By leveraging the autoregressive nature of LLMs, DISC efficiently locates difficult steps through binary partitioning, focusing compute on challenging regions rather than wasting resources on trivial steps.
-DISC is generalizable and requires no human supervision, domain-specific heuristics, prompt engineering, or process annotations, making it widely applicable across tasks.
+By leveraging the autoregressive nature of LLMs, DISC efficiently locates difficult steps through dynamically proposing step sizes, focusing compute on challenging regions rather than wasting resources on trivial steps.
+DISC is generalizable, requires no human supervision, domain-specific heuristics, prompt engineering, or process annotations, and is easily parallelizable, making it widely applicable across tasks.
+Furthermore, DISC is plug-and-play with off-the-shelf search algorithms and can be naturally integrated with greedy, beam, or Monte Carlo Tree Search.
 Our main contributions are:
-  • We introduce DISC, a method for recursive partitioning and decomposing solutions during inference without human supervision, domain-specific heuristics, or process reward models.
-  • We demonstrate how DISC integrates decomposition with inference-time search, allocating compute to high-impact, difficult steps.
-  • We show that DISC improves inference scaling in terms of both sample efficiency and token efficiency.
-  • We provide insights into how LLMs reason and plan by identifying critical steps in their generation process.
+  • We introduce DISC, a method for dynamically adjusting step sizes and decomposing solutions during inference without human supervision, domain-specific heuristics, or process reward models.
+  • We demonstrate how DISC integrates decomposition directly into inference-time search, allocating compute more effectively toward high-potential solution prefixes.
+  • We show that DISC improves inference scaling in terms of both sample efficiency, token efficiency, and runtime, achieving up to 10% reduction in error relative to the baselines and up to 4x increase in accuracy over the base model, with just 10 samples, including with reasoning models.
+  • We provide both empirical and theoretical insights into LLM reasoning, including identifying critical intermediate steps and analysis of how DISC helps discover optimal solutions.
 2 Preliminaries
-Figure 1: Comparison of different automatic decomposition methods based on step size determination.
+Figure 1: Comparison of automatic decomposition strategies based on step size.
+Coarser steps accelerate the search process but risk skipping over optimal solutions and committing to suboptimal prefixes.
+In contrast, finer steps ensure more precise decisions but lead to slower search.
+A dynamic strategy that adapts step size based on LLM feedback offers a balanced approach, combining the efficiency of coarse steps with the precision of fine-grained decomposition.
 2.1 Problem Setting
-We consider a reasoning and code generation setting where a dataset $\mathcal{X}=\{\boldsymbol{x}^{(i)}\}_{i=1}^{N}$ consists of problem prompts $\boldsymbol{x}$, and a reward model $R:\mathcal{X}\cdot\mathcal{Y}\rightarrow[0,1]$ evaluates generated solutions $\boldsymbol{y}\in\mathcal{Y}$.
-This includes program synthesis, where correctness is verified using ground-truth tests (Chen et al., 2021; Austin et al., 2021), and mathematical reasoning, where solutions are validated numerically (Hendrycks et al., 2021a; Cobbe et al., 2021).
-The reward model can be a ground-truth verifier, a trained heuristic (Zhang et al., 2024), self-consistency (Wang et al., 2023a), or an LLM-as-a-judge (Zheng et al., 2023).
-Since our focus is on step decomposition rather than verification, we use the ground-truth reward model where available.
-We assume access to a pretrained language model $\pi$ that generates text autoregressively.
-A generated response $\boldsymbol{y}$ consists of both the final solution and the reasoning chain leading to it, and can be represented as a sequence of tokens $\boldsymbol{y}=(y_{0},...,y_{L_{\boldsymbol{y}}})$.
-Additionally, solutions can be partitioned into solution steps $\boldsymbol{y}=(\boldsymbol{y}_{0},...,\boldsymbol{y}_{K})$, where each step $\boldsymbol{y}_{i}$ is a contiguous string of tokens.
-A partial solution up to step $k$ is defined as $\boldsymbol{y}_{1...k}:=\boldsymbol{y}_{1}\cdot\boldsymbol{y}_{2}\cdot...\cdot\boldsymbol{y}_{k}$, and its rollout or completion, denoted $\boldsymbol{y}_{1...k+}$, is the continuation generated by $\pi$ until an end-of-sequence token (EOS).
-The size of a solution step, $|\boldsymbol{y}_{i}|$, refers to its length in tokens or characters.
-2.2 Prior Automatic Decomposition Methods
+We consider a reasoning and code generation setting where we are given: a dataset $\mathcal{X}=\{\boldsymbol{x}^{(i)}\}_{i=1}^{N}$, a pretrained, autoregressive LLM $\pi$ that generates solutions $\boldsymbol{y}\in\mathcal{Y}$, and a reward model $R:\mathcal{X}\cdot\mathcal{Y}\rightarrow[0,1]$ that evaluates the generated solutions.
+The goal is to find inputs to the LLM that produce solutions with high reward.
+This setting includes program synthesis, where correctness is verified using ground-truth tests (chen2021evaluating, ; austin2021program, ), and mathematical reasoning, where solutions are validated numerically (hendrycks2021measuring, ; cobbe2021training, ).
+The reward model can be a ground-truth verifier, a trained heuristic (zhang2024generativeverifiersrewardmodeling, ), self-consistency (wang2023selfconsistency, ), or an LLM-as-a-judge (zheng2023judgingllmasajudgemtbenchchatbot, ) and because our focus is on decomposition rather than verification, we use the ground-truth reward model where available.
+We use the following hierarchical token notation: $y$ is a token, $\boldsymbol{p}$ is a prefix that starts at the prompt and concatenates multiple steps, $\boldsymbol{s}$ is a suffix that concatenates multiple steps and ends with the EOS token, and $\boldsymbol{y}$ is a complete solution that starts at the prompt, concatenates multiple steps, and ends at the end of sequence token EOS.
+We denote a concatenation of two tokens or token sequences with $\cdot$, e.g.
+$\boldsymbol{p}\cdot\boldsymbol{s}$ is the concatenation of prefix $\boldsymbol{p}$ and suffix $\boldsymbol{s}$ to form a complete solution $\boldsymbol{y}$.
+We denote the sampled suffix from prefix $\boldsymbol{p}$ using the LLM policy as: $\boldsymbol{s}\sim\pi(\cdot|\boldsymbol{p})$.
+We denote an optimal solution with $\boldsymbol{y}^{*}$ and an optimal suffix as $\boldsymbol{s}^{*}$, where there exist multiple optimal solutions.
+For our analysis, we use the convention that $\pi(\boldsymbol{s}^{*}|\boldsymbol{p})=0$ if there does not exist a completion $\boldsymbol{s}^{*}$ such that $\boldsymbol{p}\cdot\boldsymbol{s}^{*}$ is optimal.
+The size of a string $|\boldsymbol{y}|$, refers to its length in tokens or characters.
+2.2 Existing Decomposition Methods
 Single-step generation.
 In a single-step generation, the entire solution is generated in one pass from the prompt to the EOS token, treating it as a single action.
-This approach underlies the widely used inference scaling method best of n (BoN) (Cobbe et al., 2021; Lightman et al., 2023; Snell et al., 2024; Liang et al., 2024), where $n$ complete solutions are sampled, and the highest-scoring one is selected.
-Single-step generation also plays a role in alignment and fine-tuning methods such as DPO (Rafailov et al., 2024) and RLOO (Ahmadian et al., 2024).
+This approach underlies the widely used inference scaling method best of n (BoN) (cobbe2021training, ; lightman2023let, ; snell2024scaling, ; liang2024improving, ), where $n$ complete solutions are sampled, and the highest-scoring one is selected.
+Single-step generation also plays a role in alignment and fine-tuning methods such as DPO (rafailov2024direct, ) and RLOO (ahmadian2024back, ).
 Token-level decomposition.
 At the opposite end of the spectrum, token-level decomposition treats each atomic token as an individual step.
-While this approach dramatically increases search complexity, it enables fine-grained search that can yield higher performance gains given sufficient compute (Feng et al., 2023).
+While this approach dramatically increases search complexity, it enables fine-grained search that can yield higher performance gains given sufficient compute (feng2023alphazero, ).
 Newline and sentence-level decomposition.
-A commonly used decomposition method segments LLM generations into sentences or lines based on delimiters such as periods or newlines (Hao et al., 2023; Feng et al., 2023; Yao et al., 2024).
+A commonly used decomposition method segments LLM generations into sentences or lines based on delimiters such as periods or newlines (hao2023reasoning, ; feng2023alphazero, ; yao2024tree, ).
 Typically, each newline corresponds to a new paragraph, equation, or line of code, which often encapsulates a distinct reasoning step.
-Problem: Automatic and scalable decomposition Existing decomposition methods are task-specific, manual, and static, limiting their adaptability and scalability.
-2.3 Step Sampling
-Inference-time scaling methods must balance exploration at the current step with exploration of future steps.
-We implement a simple dynamic sampling process, referred to as negative binomial sampling, where we continue sampling completions until the sum of their rewards exceeds a predefined threshold $\sigma$.
-More formally, the number of samples $M$ drawn from a partial solution $\boldsymbol{y}_{1...k}$ is the smallest integer satisfying $\sum_{i=1}^{M}R(\boldsymbol{x}\cdot\boldsymbol{y}_{1...k+}^{(i)})\geq\sigma$.
-Where $\boldsymbol{y}_{1...k+}^{(i)}$ represents the $i$-th sampled completion from the partial solution.
-This process ensures efficient allocation of compute by dynamically adjusting the number of samples per step.
-It achieves this by either: (a) continuing to sample until a sufficiently high-reward completion is found, or (b) stopping early when additional sampling is unlikely to yield significant improvements, thus redirecting the compute to future steps.
-The stopping criterion is governed by $\sigma$: when accumulated reward from completions surpasses $\sigma$, the method assumes further sampling is unnecessary.
-For a fair comparison, we apply this sampling method uniformly across all decomposition methods in our experiments.
-2.4 Inference Scaling Methods and Decomposition
-Since our study focuses on decomposition rather than search, we primarily use greedy step search as the search method.
-In greedy step search, multiple candidate steps are sampled at each iteration, but only the highest-scoring step is retained, while the rest are discarded.
-The process then repeats, conditioning future steps on the best step found so far.
-We also perform ablation studies comparing Monte Carlo Tree Search (MCTS)  (Feng et al., 2023; Light et al., 2024b) and beam search  (Xie et al., 2024), two commonly used inference scaling methods.
-These comparisons, presented in Sec.
-4.2, highlight how different search strategies interact with decomposition.
-Additional details on MCTS and beam search are provided in App.
-E.
+Challenge: Automatic and scalable decomposition Existing decomposition methods are static and manually designed, resulting in either slow convergence to good performance or fast convergence to poor performance.
+We propose adaptive decomposition for fast convergence to a good performance.
+2.3 Search for Inference Scaling
+We differentiate between decomposition and search, where decomposition controls the number of steps between new branches and search is the process of selecting which branch to explore.
+In other words, decomposition is the construction of nodes and edges, and search is a process that occurs on that structure.
+In this work, we propose a decomposition method that is plug-and-play with different search methods.
+In our experimental results in Sec.
+4.1, we compare decomposition methods: token-level decomposition, sentence-level decomposition, and DISC with the search methods: greedy, beam (xie2024self, ), and Monte Carlo Tree Search (feng2023alphazero, ; light2024scattered, ).
+Implementation details are provided in App.
+F.
+Figure 2: Multiple iterations of Alg. 1. DISC dynamically refines its step sizes across iterations, advancing and contracting the prefix at which it samples from.
 3 Methodology
-Figure 2: DISC recursively partitions the most challenging step – i.e., the one with the lowest priority metric $h$ – to progressively refine and find the optimal solution.
-3.1 DISC Algorithm
-The DISC algorithm employs recursive binary decomposition to iteratively break down complex solutions into smaller, more manageable steps.
-Given a problem prompt $\boldsymbol{x}$, the algorithm outputs a decomposition of a solution, $\boldsymbol{y}=(\boldsymbol{y}_{1},\boldsymbol{y}_{2},\dots,\boldsymbol{y}_{K})$, such that the concatenation $\boldsymbol{y}_{1...K}$ forms a complete solution to $\boldsymbol{x}$.
-The algorithm operates in two key stages:
+The DISC algorithm uses LLM rollout data to dynamically decompose reasoning steps, adjust step sizes, and allocate sampling compute.
+In Sec.
+3.2, we present DISC paired with a greedy search strategy, shown in Alg.
+1 and Fig.
+3.
+In Sec.
+3.4, we present the general case of DISC as a nodal expansion operator which is used for pairing with beam search and MCTS.
+3.1 High-level Overview
+At a high level, DISC with Greedy Search advances a base prefix $\boldsymbol{p}_{b}$ forward by iteratively concatenating promising steps to it.
+The core intuition is to dynamically allocate compute by adjusting the step size: if a prefix shows promising reward improvement, we take a large step forward; if not, we contract the step and concentrate sampling around that prefix.
+This adaptive behavior focuses the LLM’s effort on regions of the search space that are more likely to yield high-reward completions.
+Fig.3 illustrates a single iteration of this decision process, where a candidate prefix is either accepted and extended or rejected and contracted.
+Over multiple iterations, this process yields a full solution composed of several such accepted steps.
+Fig.2 shows how the prefix is incrementally constructed: it displays the number of steps DISC has committed to the prefix, and how the prefix used for sampling dynamically changes over time.
+Together, these figures highlight the local step-wise decision-making and the global trajectory of prefix refinement throughout the search.
+Assumptions.
+DISC makes minimal, broadly applicable assumptions, enabling generality and ease of deployment.
+It avoids handcrafted prompts, process-level reward models, and domain-specific heuristics.
+Its only requirement is access to a scalar outcome reward model (ORM) to guide search.
+In the absence of a ground-truth ORM, self-supervised signals—like LLM critiques or unit tests—serve as effective substitutes (chen2022codet, ; mcaleese2024llm, ; gu2024survey, ).
+It also assumes the underlying policy $\pi$ can generate continuations from any prefix $\boldsymbol{p}$, a standard feature of decoder-only language models.
+Figure 3: DISC with Greedy Search.
+One iteration of Alg.
 1.
-Solution sampling.
-Starting from $\boldsymbol{x}$ and $\boldsymbol{y}_{0}=\emptyset$, the algorithm generates complete solutions $\boldsymbol{y}_{k}\sim\pi(\cdot|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(1-k)})$ using a policy $\pi$ like in single step generation.
-The best solution, $\boldsymbol{y}^{*}_{k}$, is selected based on the reward model $R$.
-2.
-Recursive partitioning.
-The selected solution $\boldsymbol{y}^{*}_{k}$ is partitioned into two segments, $\boldsymbol{y}^{*}_{k}=\boldsymbol{y}^{*}_{a}\cdot\boldsymbol{y}^{*}_{b}$, based on a predefined partition fraction $\alpha$, where $|\boldsymbol{y}^{*}_{a}|\approx\alpha|\boldsymbol{y}^{*}_{k}|$.
-For each part, a priority metric $h$ is estimated: $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ and $\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, usually through rollouts of the step using $\pi$.
-The part with the lower priority is further partitioned.
-  • If $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})\geq\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, additional samples are sampled for $\boldsymbol{y}^{*}_{b}$, with the process repeating on the new best solution, $\boldsymbol{y}^{\prime*}_{b}$.
-  • Conversely, if $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})<\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, the first segment $\boldsymbol{y}^{*}_{a}$ is further partitioned.
-The first step corresponds to the $\alpha$ fraction of $\boldsymbol{y}^{\prime*}_{a}$, with the remaining part of the full solution forming the second step.
-This recursive process is illustrated in Fig. 2. The pseudocode for DISC is provided in Alg. 1, with an annotated Python implementation in App. A.
-The priority metric $h$ serves as the central heuristic for determining which solution steps to prioritize.
-It estimates the “difficulty” or “potential for improvement” of a step $\boldsymbol{y}_{k}$ given the context $\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)}$, computed via rollouts of the policy $\pi$.
-Specifically, $h(\boldsymbol{y}_{k}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ is estimated by sampling continuations and evaluating their outcomes.
-In practice, estimating $h$ and generating new samples occur simultaneously, as both rely on rollout-based computations (Sec.
-3.2).
-Unlike standard decomposition methods, DISC does not process steps in strict temporal order, resembling goal-directed planning (Parascandolo et al., 2020) and backtracking.
-Intuition and Benefits.
-DISC partitions difficult steps into smaller, simpler sub-steps, allocating additional resources to refine them.
-This approach is particularly effective for inference scaling, where computing must be used judiciously.
-A binary decomposition strategy enables fast identification of difficult or high-potential steps.
-Key Insight: Recursive partitioning Top-down, recursive partitioning means that we can both efficiently locate critical steps and also dynamically determine step sizes based on our budget.
-Dynamic Compute Allocation.
-A key advantage of DISC is its ability to prioritize challenging or high-potential steps, improving solution quality while minimizing compute waste.
-By iteratively refining steps with low priority scores, DISC adaptively allocates more resources to difficult steps and less to simpler ones, optimizing inference efficiency.
-Key Insight: Adaptive compute allocation DISC dynamically allocates inference compute to harder steps, optimizing solution quality and resource efficiency.
-3.2 Priority Metric
-We consider two intuitive priority metrics for step selection: Q-value priority and Z-score priority, which are visualized in Fig. 4.
-Q-value based priority (DISC-Q).
-Given a partial solution $\boldsymbol{y}^{*}_{1...k}=\boldsymbol{y}^{*}_{1...(k-1)}\cdot\boldsymbol{y}^{*}_{a}\cdot\boldsymbol{y}^{*}_{b}$, we aim to prioritize either $\boldsymbol{y}^{*}_{a}$ or $\boldsymbol{y}^{*}_{b}$ for refinement.
-The core intuition behind Q-value prioritization is that steps with lower Q-values indicate areas needing refinement, directing compute toward the most challenging parts of the solution.
-More formally, we define the Q-priority metric under policy $\pi$ as:
-$h_{Q}(\boldsymbol{y}^{*}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})=\mathbb{E}_{\boldsymbol{y}_{k}}\left[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})\right].$
-Here, $\boldsymbol{y}_{k}$ represents alternative steps sampled from $\pi$, and $Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ denotes the Q-value of $\boldsymbol{y}_{k}$, conditioned on the partial solution $\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)}$.
-Equivalently, $h_{Q}(\boldsymbol{y}^{*}_{k})$ can be interpreted as the value function $V^{\pi}(\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$, where $V^{\pi}$ represents the expected reward achievable from the given partial solution.
-The expectation is taken over sampled candidates $\boldsymbol{y}_{k}\sim\pi(\cdot\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$, constrained by $|\boldsymbol{y}_{k}|=|\boldsymbol{y}^{*}_{k}|$.
-This metric helps identify difficult steps that the LLM is likely to get wrong, guiding partitioning toward the most critical refinements.
-To estimate $h_{Q}$ for the second step $\boldsymbol{y}^{*}_{b}$, we sample $\boldsymbol{y}_{b}$ from $\pi$, generate rollouts $\boldsymbol{y}_{(k+1)+}$, compute rewards, and average the outcomes.
-For the previous step $\boldsymbol{y}^{*}_{a}$, we reuse rollouts from earlier partitioning, as the mean of these previously generated samples provides an unbiased estimate of the Q-priority metric:
-$h_{Q}(\boldsymbol{y}^{*}_{a}\mid\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...(k-1)})=\mathbb{E}\left[R(\boldsymbol{y}^{*}_{1...(k-1)+})\right].$
-By leveraging existing rollouts, DISC avoids redundant sampling, improving computational efficiency.
-Once the rollouts for $\boldsymbol{y}^{*}_{(k+1)}$ are available, the best completion $\boldsymbol{y}^{*}_{(k+1)+}$ is selected as the next step to partition.
-This dual use of rollouts optimizes both metric estimation and inference sampling.
-Combining priority estimation and test-time search By integrating LLM policy sampling for both metric estimation and search, we can significantly enhance computational efficiency.
-Figure 3: Priority metric estimation. We can estimate the priority metric $\hat{h}$ using Monte Carlo rollouts of the LLM policy $\pi$ for each step.
-Z-score based priority (DISC-Z).
-To allocate more compute to steps with higher potential for improvement, we estimate the probability of sampling a better step given existing samples.
-Assuming the Q-values of sampled steps follow a normal distribution, we model this probability using the cumulative distribution function (CDF).
-Given a mean $\mu_{r}$ and standard deviation $\sigma_{r}$ of sampled Q-values, the probability of sampling better than the best-observed step $\boldsymbol{y}^{*}_{k}$ with Q-value $q^{*}$ is:
-$1-\operatorname{CDF}\left(\frac{q^{*}-\mu_{r}}{\sigma_{r}}\right)=1-\operatorname{CDF}(z^{*}),$
-where $z^{*}$ is the Z-score. Since the CDF is monotonic, we compare steps based on their Z-scores.
-We formally define the Z-score priority metric as:
-$h_{Z}(\boldsymbol{y}^{*}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})=\frac{q^{*}-\mathbb{E}[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})]}{\operatorname{Std}[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})]}.$
-Since $q^{*}-\mu_{r}$ represents the advantage of step $\boldsymbol{y}^{*}_{k}$, the Z-score metric can be interpreted as a standard deviation-scaled advantage.
-Lower $h_{Z}$ values indicate steps with greater room for improvement, guiding decomposition toward those with higher variance in performance.
-Figure 4: Different priority metrics. Sampled rewards of different steps can be visualized as a distribution.
-3.3 DISC and Search Methods
-DISC can also be used to enhance Monte Carlo Tree Search (MCTS) and other inference scaling and search methods.
-Recall that in our partition step, we greedily partition the best solution step $\boldsymbol{y}^{*}$.
-Instead of greedily partitioning the best step, we can partition the top $k$ best steps instead, and select which step to partition using the upper confidence tree (UCT) formula.
-We can also use beam search to prune out steps we do not want to partition further.
-We explain MCTS and beam search in detail in App.
-E and present results of combining DISC with search in Sec.
-4.9.
-Takeaway: Dynamic step sizes can improve search DISC can enhance search based inference scaling methods by determining what step size to search across.
-3.4 A Motivating Example on DISC-Z
+We start with a base prefix A and a candidate prefix B.
+We compare the sample statistics of each by evaluating a scoring function (e.g., z-score).
+If the candidate prefix B demonstrates a higher likelihood of reward improvement compared to continuing from base prefix A, we accept B, commit it as the new base, and extend the candidate to a further step (e.g., BD) on the best sampled solution.
+If B is not better, we reject it and propose a shorter candidate (e.g., AC), contracting the step size.
+This process repeats until a new candidate is accepted or all options are exhausted.
+The algorithm thus adaptively advances or contracts the step size and search horizon based on the relative quality of completions from each prefix.
+3.2 DISC with Greedy Search Algorithm
+Figure 4: Reward distribution of rollouts sampled from two different prefixes. The probability of sampling a higher rollout from prefix 2 is higher than that of prefix 1.
+We now describe Alg.
+1.
+The algorithm takes as input an LLM $\pi$, a reward model $R$, prompt $\boldsymbol{x}$, initial partition fraction $\alpha_{0}$, threshold $\sigma$, and sample budget $N$.
+It initializes the base prefix as $\boldsymbol{p}_{b}=\boldsymbol{x}$ and sets $\alpha=\alpha_{0}$.
+At each iteration, the algorithm samples $\pi$ from $\boldsymbol{p}_{b}$ to generate completions $\boldsymbol{y}^{i}=\boldsymbol{p}_{b}\cdot\boldsymbol{s}^{i}$, computing rewards $R(\boldsymbol{y}^{i})$.
+Sampling stops when the cumulative reward exceeds $\sigma$: $M=\min\{m\in\mathbb{Z}_{>0}\mid\sum_{i=1}^{m}R(\boldsymbol{y}^{i})\geq\sigma\}$.
+This balances sample quantity and quality.
+The solutions and their rewards are stored in $Y$, and the best suffix $\boldsymbol{s}^{*}$ and z-score are computed.
+A candidate prefix $\boldsymbol{p}_{c}$ is then formed by appending the first $\alpha$ fraction (token-wise) of $\boldsymbol{s}^{*}$ to $\boldsymbol{p}_{b}$.
+It is accepted if its reward z-score $z_{c}$ is lower than the current $z_{b}$.
+Assuming rewards follow a location-scale family distribution (e.g., Gaussian), a lower z-score implies a higher probability of improvement, since $\Pr(R>\text{max}_{\boldsymbol{s}\in S_{b,c}}(R(\boldsymbol{s}))=1-\mathrm{CDF}(z_{b,c})$ where $\mathrm{CDF}$ is the cumulative distribution function.
+Fig.
+4 displays how we can estimate the probability of improvement from a reward distribution.
+A location-scale distribution of rewards is supported empirically (see App.
+C.5).
+If the candidate is accepted, the algorithm updates the base prefix $\boldsymbol{p}_{b}\leftarrow\boldsymbol{p}_{c}$, resets $\alpha=\alpha_{0}$, and updates the base z-score and best suffix.
+If rejected, the partition fraction contracts: $\alpha\leftarrow\alpha\cdot\alpha_{0}$.
+This contraction implements DISC’s recursive refinement mechanism, focusing the search on more promising regions.
+The process repeats until the sample budget is exhausted or a correct solution is found.
+Algorithm 1 DISC with Greedy Search
+1:LLM $\pi$, Reward model $R$, prompt $x$, initial partition fraction $\alpha_{0}$, negative binomial threshold $\sigma$, sample budget NN
+2: // initialize base prefix and current partition fraction
+[…]
+4: // compute base prefix statistics (_b)
+[…]
+[…]
+[…]
+8:while $n<N$ do
+9:   // get candidate prefix (_c)
+10:  $\boldsymbol{p}_{c}=\boldsymbol{p}_{b}\cdot\textup{split}(\boldsymbol{s}_{b}^{*},\alpha)$
+11:   // sample and compute candidate prefix statistics (_c)
+12:  $S_{c}=\{\boldsymbol{p}_{c}\cdot\boldsymbol{s}^{i}\mid\boldsymbol{s}^{i}\sim\pi(\cdot|\boldsymbol{p}_{c})\}_{i=1}^{M}$
+13:  $z_{c}=\frac{\text{max}_{\boldsymbol{s}\in S_{c}}(R(\boldsymbol{s}))-\text{mean}_{\boldsymbol{s}\in S_{c}}(R(\boldsymbol{s}))}{\text{std}_{\boldsymbol{s}\in S_{c}}(R(\boldsymbol{s}))}$
+14:  $\boldsymbol{p}_{c}\cdot\boldsymbol{s}^{*}_{c}=\text{argmax}_{\boldsymbol{s}\in S_{c}}R(\boldsymbol{s})$
+15:  $n\leftarrow n+M$
+16:   // accept or reject the candidate prefix
+17:  if $z_{c}<z_{b}$ or $|\boldsymbol{s}_{b}^{*}|\leq 1$ then
+18:   $\boldsymbol{p}_{b}\leftarrow\boldsymbol{p}_{c}$, $\boldsymbol{s}^{*}_{b}\leftarrow\boldsymbol{s}^{*}_{c}$,
+19:   $\alpha\leftarrow\alpha_{0}$, $z_{b}\leftarrow z_{c}$
+20:  else
+21:   $\alpha\leftarrow\alpha_{0}\alpha$
+22:  end if
+23:end while
+24:yield $\boldsymbol{p}_{b}\cdot\boldsymbol{s}^{*}_{b}$
+3.3 Analysis of DISC with Greedy Search
+DISC exhibits two important properties: (i) the z-score decreases monotonically over the course of algorithm iterations, and (ii) the best candidate solution always has a higher reward than the best base solution.
+We leverage these properties, together with assumptions about the quality of $\pi$, to establish the following result.
+We also develop a motivating theoretical example in App.
+G.2.
+Theorem 1 (Optimality of DISC)
+Consider Alg.
+1.
+Suppose that for some problem $\boldsymbol{x}$, the optimal solution is in the support of $\pi(\cdot\mid\boldsymbol{x})$.
+Then at some $n>0$, the base prefix contains EOS token, the algorithm terminates, and this solution is an optimal solution.
+See App.
+G for proof.
+Remark 1
+Our analysis is dependent on a strong policy assumption, which is "reverse engineered" to be the weakest assumption on the policy such that our algorithm terminates at optimality.
+In other words, this assumption depends on an instance-dependent property that must be checked empirically.
+The purpose of our analysis is not to provide a universal guarantee, but rather to understand how adaptive decomposition method can control real-time inference compute without sacrificing optimality.
+3.4 DISC for Plug-and-Play with Search Algorithms
+In the previous section, we presented DISC with greedy search as a complete algorithm.
+However, the core of DISC is the decomposition that controls from which token prefixes to query the model and dynamically adjusts step sizes, which can be plug-and-played with other search algorithms like Beam search and MCTS.
+The DISC ExpandNode and MakeChildren are presented in Alg.
+3.
+These methods are conceptually similar to the DISC with greedy search, except that in the general search context, each node stores its own base prefix, and so when a node is expanded it computes its candidate prefix from the suffix to that node instead of from the best existing suffix.
+During the expansion operator, multiple sets of children are generated, but only the children from the final prefix are kept for search.
+3.5 Example Decomposition
+Using a sampling budget of $N=100$ LLM calls, the decomposition of a representative MATH problem is shown below.
+Each step is enclosed in brackets and color-coded based on the z-score of sampling a better suffix conditioned on the current prefix.
+Brighter colors indicate lower z-scores, signaling higher importance.
+DISC example decomposition [ Let the length of the rectangle be $l$ ] [ and the width of the rectangle be $w$.
+Since the perimeter of the rectangle is 24 inches, we have that $2l+2w=24$, so $l+w=12$.
+We wish to maximize the area of the rectangle, ] [ which ] [ is $A=lw$.
+Let l=12wl=12-w and plug into the area: $\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}A=(12-w)w\Rightarrow\qquad A=12w-w^{2}$} Now, we differentiate AA with respect to $w$: $\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}A^{\prime}(w)=12-2w$} We wish to maximize AA, so we set $A^{\prime}(w)=0$, and solve for $w$: $\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}12-2w=0\Rightarrow\qquad w=6$} Since l=12wl=12-w, we have that l=126=6l=12-6=6.
+Therefore, the area of the rectangle is $A=lw=6\cdot 6=\boxed{36}$.
+]
+DISC identifies the first step as highly important, which aligns with intuition—early reasoning forms the foundation for all subsequent steps.
+In contrast, the final step, although large, is marked as low-importance, indicating that DISC allocated minimal compute toward refining it.
+This suggests that once earlier reasoning is fixed, there is limited opportunity for improvement in the final conclusion using additional sampling.
+Interestingly, the third step, beginning with "which", is assigned high importance and receives substantial attention from DISC.
+This step appears to act as a pivotal decision point that shapes the direction of the remaining solution.
+This observation supports the idea that certain tokens or sub-sequences function as critical reasoning forks or pivots—consistent with findings from prior work (bigelow2024forking, ; muennighoff2025s1, ).
+Autoregressive models require autoregressive decomposition While transition words such as ‘which’, ‘therefore’, ‘wait’, etc.
+may not appear significant to human readers, our decomposition frequently identifies them as critical decision points for autoregressive LLMs trained on next-token prediction, where selecting a different token at these junctures can substantially alter the downstream reasoning and final outcome.
+Therefore, it is essential for inference algorithms to allocate more compute toward sampling at these steps, and to identify these steps automatically through LLM data rather than through human design.
+4 Experimental Results
+4.1 Main Results
+Benchmarks.
+We evaluate DISC on three benchmarks: APPS, MATH, and LiveCodeBench, to assess its impact on inference scaling for both coding and reasoning.
+APPS (hendrycks2021measuring, ) consists of 5000 competitive programming problems across three difficulty levels, with the competition-level subset being the hardest.
+We evaluate on a 200-problem subset due to computational constraints.
+MATH (hendrycks2021measuring-MATH, ) comprises 12,500 math problems.
+Since the ground-truth verifier provides only binary rewards, we use a pretrained ORM (xiong2024rlhflowmath, ), trained via the method in (wang2024math, ), with Llama-3.1-8B-Instruct as the base model.
+We test on a 500-problem subset (MATH500), identical to prior work (wang2024math, ; lightman2023let, ).
+LiveCodeBench (jain2024livecodebench, ) is a continuously updated dataset from Leetcode, AtCoder, and CodeForces, ensuring LLMs have not been exposed to test problems.
+We evaluate on the 108 problems uploaded between 10/01/2024 and 12/01/2024 to prevent contamination.
+Figure 5: Token-level comparisons across benchmarks using gpt-4o-mini.
+(Left) APPS competition level (Middle) MATH500 (Right) LiveCodeBench.
+DISC achieves superior inference scaling over baselines on all three benchmarks.
+Baselines.
+We compare DISC against three prior decomposition methods: TokenSplit (token-level decomposition), LineSplit (newline-based decomposition), and BoN (treating the entire solution as a single step).
+These are all standard and the most commonly used methods (see Sec.
+2.2).
+Metrics.
+We evaluate two key metrics: Pass@k, the proportion of problems solved within a sample budget $k$, and Pass@token, the proportion solved within a given token budget.
+Note that $k$ refers to the sample budget, not thousands of samples, and error proportion refers to proportion not solved.
+We use $\alpha_{0}=0.15,\sigma=1.0$, and temperature $\tau=0.2$ by default for DISC unless otherwise specified.
+Performance.
+Across all benchmarks, DISC consistently delivers stronger performance and better scaling under both fixed token budgets (Fig.
+5) and fixed sample budgets (Fig.
+8).
+For example, on APPS, the pass@10 error proportion decreases from 0.50 to 0.475; on MATH500, from 0.15 to 0.14; and on LiveCodeBench, from 0.57 to 0.51.
+These correspond to a 5.0%, 6.7%, and 10.5% reduction in error relative to the best baseline, respectively.
+These improvements are particularly meaningful on more challenging benchmarks, where performance gains are harder to achieve, demonstrating DISC ’s effectiveness in guiding search toward high-reward regions.
+Extended results and analyses are provided in App.
+D.1D.4, and D.5.
+Figure 6: Inference scaling for different models, including reasoning models, on APPS.
+(Left Pass@k for open-source models.
+Colors correspond to different models, with Pass@1 being the base model performance.
+(Middle) Pass@k for R1, with green line as base Pass@1 performance.
+(Right) Pass@token for R1.
+DISC almost doubles base performance across all models.
+Model Generality.
+DISC also provides substantial gains across a range of models, including open-source LLMs.
+For instance, it improves LLaMA’s pass@10 rate from 0.01 to 0.04—a 300% relative increase.
+Similarly, it boosts Mistral’s performance from 0.0 to 0.02, and Qwen’s from 0.095 to 0.17, representing a 79% relative increase.
+As shown in Fig.
+6 (left), these improvements hold consistently across sampling budgets, including in low-budget regimes.
+This demonstrates DISC ’s general applicability and effectiveness even for weaker or resource-constrained models.
+Additional results and analyses are provided in App.
+C.3.
+Reasoning Models.
+DISC also yields substantial gains when applied to strong long-form chain-of-thought (CoT) reasoning models such as R1 (guo2025deepseek, ; chen2025towards, ).
+As shown in Fig.
+6 (middle), DISC improves accuracy by over 85% relative to the base R1 model using just 10 samples.
+Notably, Fig.
+6 (right) shows that even under a constrained token budget—matched to that of a single sample from the base model—DISC still achieves over a 33% relative improvement.
+This demonstrates that DISC not only scales well with more samples but is also highly effective at identifying and prioritizing critical reasoning steps, leading to stronger performance even in low-resource settings.
+Computational Overhead.
+DISC introduces negligible runtime overhead compared to standard decoding baselines, as shown in Fig.
+7.
+The vast majority of compute time—over 90% across all settings—is still dominated by LLM token sampling, with only a minor fraction spent on auxiliary operations such as candidate management, z-score normalization, and recursive branching.
+Despite its dynamic control flow, DISC maintains a runtime composition nearly identical to that of methods like BoN and LineSplit.
+Moreover, because DISC achieves higher success rates with fewer tokens (Fig.
+5), its effective runtime per correct solution is even lower.
+In practice, this makes DISC both algorithmically efficient and computationally scalable, offering improved search performance without additional inference cost.
+We provide a detailed runtime discussion in App.
+D.6.
+Figure 7: Percentage of runtime spent on overhead vs LLM token generation. DISC does not increase the runtime overhead significantly.
+Figure 8: Comparison of Pass@k performance on APPS using gpt-4o-mini. (Left) self generated validation tests, (Middle) with ground truth tests, (Right) with different search methods.
+Self-Generated Validation.
+We further evaluate DISC in a self-generated validation setting, where ground-truth reward models or curated tests are unavailable (chen2022codet, ; chen2023teaching, ; zhou2023language, ).
+Here, the LLM generates its own unit tests from the problem description, which serve as a proxy reward model for evaluating candidate solutions.
+This setup offers a scalable alternative to costly manual test curation in real-world code generation tasks.
+As shown in Fig.
+8, DISC scales effectively under this protocol, achieving a 54% relative improvement over the base model.
+This demonstrates that DISC leverages decomposition-based reasoning to produce higher-quality code even when supervision is noisy or incomplete.
+However, self-generated validation has limitations: generated tests may be incomplete, inconsistent, or narrow in coverage, potentially biasing performance estimates.
+While these issues do not affect the observed scaling trends, improving test reliability and robustness remains an important direction.
+Additional details and results are provided in App.
+D.3.
+Search.
+We demonstrate that search strategies such as MCTS and beam search can be naturally integrated with DISC using the approach described in Sec.
+3.4.
+As shown in Fig.
+47, greedy search tends to explore deeper partitions within the same search budget due to its myopic nature, while MCTS and beam search explore more diverse but shallower paths.
+Despite similar depth, MCTS outperforms beam search by allocating its search budget more strategically—focusing exploration on more promising candidates—resulting in superior overall performance, as seen in Fig.
+8.
+Furthermore, unlike greedy search, which irrevocably commits to prefixes as they are accepted, MCTS maintains flexibility by exploring committing multiple candidate prefixes in parallel.
+Additional details and analysis are provided in App.
+F.
+4.2 Ablation Studies
+Temperature.
+Figure 9: Ablation study of DISC on APPS with gpt-4o-mini.
+(Left) Effect of temperature: unlike BoN and other inference scaling methods, DISC achieves higher performance at lower temperatures.
+(Middle) effect of acceptance method (Right).
+Effect of partition fraction $\alpha_{0}$: The range $0.15\leq\alpha_{0}\leq 0.25$ appears optimal.
+Typically, inference scaling methods achieve optimal performance at temperatures around 0.6–0.8, as increased temperature promotes sample diversity (wang2024planning, ).
+Surprisingly, however, DISC performs better at lower temperatures, as shown in Fig.
+9.
+This trend is in stark contrast to BoN (Fig.
+16), where higher temperatures are generally beneficial.
+We believe this phenomenon arises because DISC depends on estimating the z-score at each step using sample statistics.
+Lower temperatures reduce sample variance, leading to more reliable estimates, which in turn improves step selection.
+This is further supported by Fig.
+16, which shows that lower temperatures yield lower standard deviations per step, indicating increased sampling consistency.
+Additional details and analyses can be found in App.
+C.1.
+Acceptance Method.
+We perform an ablation study to evaluate whether using the z-score is an effective criterion for accepting candidate prefixes.
+Specifically, we compare our standard z-score-based acceptance method, DISC-Z, against four alternative baselines: DISC-R, which accepts candidates uniformly at random; DISC-Q, which accepts if the candidate prefix has a lower mean value; DISC-negQ, which accepts if the candidate has a higher mean; and DISC-negZ, which accepts if the candidate has a higher z-score (rather than a lower one).
+As shown in Fig.
+9, the choice of acceptance criterion substantially influences performance.
+Among all methods, DISC-Z achieves the highest performance, while DISC-negZ performs worse than random selection, underscoring the importance of prioritizing candidates with a higher probability of improvement.
+Additional details and analysis are in App.
+C.2.
+Partition Fraction $\alpha_{0}$.
+As shown in Fig.
+9 and Fig.
+26, performance is highest when the partition fraction lies in the range $0.15\leq\alpha_{0}\leq 0.25$.
+Smaller values of $\alpha_{0}$ generally yield better results because they lead to more conservative proposals—i.e., shorter candidate prefixes.
+This conservatism is beneficial due to the high cost of prematurely committing to a suboptimal prefix: once a candidate prefix is accepted (in greedy-DISC), it becomes fixed and cannot be revised.
+By keeping candidate prefixes short, the algorithm retains more flexibility to correct course in future steps.
+Additional analysis is provided in App.
+C.4.
+4.3 Analysis and Interpretation
+Figure 10: Analysis of decomposition methods.
+Dotted lines fit a linear model to indicate the trend.
+(Left) Average reward per step: From step 3 onward, higher step counts strongly correlate with increased average reward, demonstrating the effectiveness of decomposition.
+The dip between steps 1 and 3 likely occurs because simple problems are solved early, preventing further search.
+(Middle) Standard deviation of rewards per step: Decomposition reduces sampling variance, improving precision at deeper search depths.
+(Right) Frequency of the number of partition steps that the search algorithm committed to the prefix during the search process.
+Our results strongly suggest that decomposition—whether line-based, token-based, or DISC —consistently improves sample quality.
+Fig.
+10 (left and middle) shows how the mean and variance of sampled rewards evolve with the step number, i.e., the number of steps committed to the prefix during search.
+As the step number increases, the mean reward improves, indicating that longer committed prefixes lead to higher-quality solutions.
+At the same time, the variance of rewards decreases, suggesting that committing to a longer prefix also improves the precision of sampling.
+These trends highlight the benefits of finer-grained decomposition and incremental commitment in guiding the search process more effectively.
+Furthermore, DISC achieves higher performance using fewer committed prefix steps—and thus fewer sampling stages—under a fixed sampling budget.
+Fig.
+10 (right) shows the distribution of actual partitions, i.e., the number of steps effectively committed under a finite budget.
+As shown, DISC typically requires only 1–5 actual partition steps, whereas other methods commit to significantly more.
+This indicates that DISC is more efficient at identifying high-impact prefixes, enabling better performance with fewer sampling decisions.
+Limitations.
+While DISC demonstrates strong empirical performance, several limitations remain (see Appendix E for a detailed discussion).
+In particular, DISC assumes access to a reasonably accurate reward model.
+For example, in code generation tasks, this requires that either ground-truth validation tests are available or the model can self-generate sufficiently reliable ones.
+Addressing these limitations—through improved test generation, learned verifiers, and more robust evaluation protocols—presents promising directions for future work.
+Decomposition and Sample Quality DISC enables more efficient exploration by identifying high-impact prefixes with fewer steps.
+Incremental prefix commitment not only improves sample quality—yielding higher average rewards—but also reduces reward variance, leading to more stable and reliable outputs under a fixed sampling budget.
+5 Related Work
+Inference scaling.
+Inference scaling has emerged as a dominant paradigm, driven by the introduction of o1- and r1-like chain-of-thought reasoning models (snell2024scaling, ; brown2024large, ; manvi2024adaptive, ; leea2025evolving, ).
+Several works examine the trade-off between inference compute and training compute (guan2025rstarmathsmallllmsmaster, ; chen2024think, ).
+LLM inference often relies on decomposing complex problems into intermediate reasoning steps, as seen in chain-of-thought (CoT) prompting (wei2022chain, ; sprague2024cotcot, ; wang2024chainofthoughtr, ) and its variants (kojima2022large, ; zhouleast, ; wangself, ; li2023making, ).
+We extend inference scaling by introducing a new approach for adaptive compute allocation (manvi2024adaptive, ; alabdulmohsin2025recursive, ; wang2025dynscaling, ).
+LLM reasoning and code generation.
+LLM reasoning and code generation are central tasks for inference scaling.
+Evolutionary inference scaling methods have been explored in program generation (liventsev2023fully, ; chen2023evoprompting, ; romera2024mathematical, ; lehman2023evolution, ; hemberg2024evolving, ).
+Domain-specific decomposition strategies have been applied in code generation, such as function-based decomposition (chen2024divide, ; zenkner2024abstractbeam, ; levin2025effective, ).
+More broadly, decomposition often involves prompting LLMs to generate subtask completions (hernandez2recursive, ; khot2022decomposed, ; dua2022successive, ), which differs from methods that refine a single LLM generation.
+Reinforcement learning and Monte Carlo methods.
+Unlike standard RL, our setting resembles a search problem where the goal is to identify the single highest-reward path.
+Nested Monte Carlo search can accelerate optimal pathfinding  (cazenave2009nested, ).
+Under the bandit setting, this can be formulated as identifying the arm with the highest maximum reward rather than the highest mean reward (cicirello2005max, ; carpentier2014extreme, ).
+6 Conclusion
+We introduce DISC, a dynamic decomposition framework that adaptively partitions solution steps based on first order statistics that capture potential for improvement, improving inference scaling by directing compute toward critical steps while balancing exploration and resource allocation.
+DISC naturally integrates with search-based methods such as MCTS and beam search, further enhancing performance.
+It also identifies challenging steps for LLMs, aiding curriculum learning, fine-tuning, and dataset augmentation.
+By dynamically adjusting partitioning and step sizes based on available compute, DISC enables more adaptive and efficient reasoning in large language models, with broad implications for both training and inference optimization.
+Appendix A Code implementation of DISC
+To aid reproducibility and practical adoption, we provide Python-style pseudocode in this section that closely mirrors our actual implementation.
+The pseudocode abstracts away low-level engineering details while preserving the core logic, function names, and control flow used in our codebase.
+This alignment ensures that readers can easily translate the pseudocode into a working implementation or modify it for their own use cases.
+All key components of the DISC algorithm—including dynamic step refinement, z-score-based acceptance, and integration with search strategies—are reflected in this pseudocode with minimal deviation from the actual code.
+Python implementation of DISC def dynamic_decomposition(problem, model, reward_model, split_str, complete_solution, fraction, solution_budget, split_metric, stop_threshold=-float("inf"), stop_sum_score=1.0, stop_if_solved=False, ): """ ␣␣␣␣Decomposesthesolutionusingadynamicbinarysearchapproach ␣␣␣␣Args: ␣␣␣␣␣␣␣␣problem(Problem):Theproblemtosolve ␣␣␣␣␣␣␣␣model(Model):Themodeltouseforgeneration ␣␣␣␣␣␣␣␣reward_model(function):Therewardmodeltouseforscoring ␣␣␣␣␣␣␣␣split_str(function):Thefunctiontouseforsplittingastring ␣␣␣␣␣␣␣␣complete_solution(function):Thefunctiontouseforcompletingasolution ␣␣␣␣␣␣␣␣fraction(float):Thefractiontosplitthestring ␣␣␣␣␣␣␣␣solution_budget(int):Themaximumnumberofsolutionstogenerate ␣␣␣␣␣␣␣␣split_metric(function):Themetrictouseforsplitting ␣␣␣␣␣␣␣␣stop_threshold(float):Thethresholdtostopsplitting ␣␣␣␣␣␣␣␣stop_sum_score(float):Thesumscoretostopgeneratingcompletions ␣␣␣␣␣␣␣␣stop_if_solved(bool):Whethertostopiftheproblemissolved ␣␣␣␣""" # Initialize results and decomposition steps decomp_return = { "generated_solutions": [], "decomposition": [] } while len(decomp_return["generated_solutions"]) < solution_budget: # Combine all previous steps into an intermediate solution intermediate_solution = "".join([step["step_str"] for step in decomp_return["decomposition"]]) new_scores = [] best_solution = None best_completion = None best_score = -float("inf") sum_score = 0.0 # 1) Generate completions until we generate enough samples to estimate the split metric while sum_score < stop_sum_score: proposed_completion = complete_solution(problem, intermediate_solution, model) proposed_solution = intermediate_solution + proposed_completion decomp_return["generated_solutions"].append(proposed_solution) # Update scores proposed_score = reward_model(proposed_solution) new_scores.append(proposed_score) sum_score += proposed_score # Track the best solution if proposed_score > best_score: best_solution = proposed_solution best_score = proposed_score best_completion = proposed_completion # Stop early if problem is solved if stop_if_solved and proposed_score >= 1.0: decomp_return["decomposition"].append({"step_str": proposed_completion}) return decomp_return new_metric = split_metric(new_scores) last_metric = decomp_return["decomposition"][-1]["metric"] if decomp_return["decomposition"] else None # Determine the split target.
+We always split the step with the highest metric is_split_new_step = last_metric is None or new_metric >= last_metric split_target = decomp_return["decomposition"][-1]["step_str"] if not is_split_new_step else best_completion # 3) Attempt to split the target split_result = split_str(split_target, fraction) if not split_result: # If we can’t split the target, we’re done decomp_return["decomposition"].append({"step_str": best_completion, "metric": new_metric}) return decomp_return # Update decomposition based on split part1, part2 = split_result if is_split_new_step: decomp_return["decomposition"].append({"step_str": part1, "metric": new_metric}) # Stopping condition based on threshold if new_metric < stop_threshold: decomp_return["decomposition"].append({"step_str": part2}) return decomp_return else: decomp_return["decomposition"][-1] = {"step_str": part1, "metric": last_metric} return decomp_return
+Appendix B Pseudocode for DISC
+Algorithm 2 Dynamic Decomposition
+[…]
+[…]
+3:Initialize ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}\leftarrow\{\text{generated\_solutions}:\varnothing,\text{decomposition}:\varnothing\}$
+[…]
+5:while $|{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{generated\_solutions}|<{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$ do
+6:  ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}}\leftarrow\text{Concatenate}([\text{step.step\_str}\ \forall\ \text{step}\in{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}])$
+[…]
+8:  $\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}$$\leftarrow\varnothing \vartriangleright$ Record rewards of completions
+9:  ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{final}}}\leftarrow\text{None}$, ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\leftarrow\text{None}$, $\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r$$\leftarrow-\infty \vartriangleright$ Track the best completion
+10:  # Step 1: Generate completions until we have enough samples to estimate the splitting metric. Here we use a geometric sampling distribution
+11:  while $\operatorname{sum}({\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}})<{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\sigma}$ do
+12:   ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\pi}(\cdot|{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}x},{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}})$
+13:   ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}}\oplus{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}$
+14:   Append $\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}$} to ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{generated\_solutions}$
+15:   ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}\leftarrow{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r}({\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}})$
+16:   Append $\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}$} to $\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}$}
+17:   if ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}>{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}$ then
+18:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{final}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}}$, ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}$
+19:     ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}\leftarrow{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}$
+20:   end if
+21:   if $\textbf{b}_{\text{inference}}\ \textbf{and}\ {\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}=1.0$ then
+22:     Append $\{\text{step\_str}:{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}\}$ to ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}$
+[…]
+24:   end if
+25:  end while
+26:  # Step 2: Compute splitting metric
+27:  ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\leftarrow{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}h}({\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}})$
+28:  ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}[-1].{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z}\ \textbf{if}\ {\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}\neq\varnothing\ \textbf{else}\$-$\infty$
+29:  # Step 3: Split the step with the higher metric
+30:  bsplit new step $\leftarrow$ ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\geq{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}$
+31:  ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{target step}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\ \textbf{if}\ \textbf{b}\textsubscript{\text{split new step}}\ \textbf{else}\ {\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}[-1].\text{step\_str}$
+32:  ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}},{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}f}({\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{target step}}},{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\alpha})$
+33:  if ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}}=\text{None}\ \textbf{or}\ {\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}=\text{None}$ then
+34:   Append $\{\text{step\_str}:{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}},\text{metric}:{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\}$ to ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}$
+[…]
+36:  end if
+37:  if $\textbf{b}_{\text{split new step}}$ then
+[…]
+39:   if ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}<{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$ then
+40:     Append $\{\text{step\_str}:{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}\}$ to ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}$
+[…]
+42:   end if
+43:  else
+[…]
+45:  end if
+46:end while
+[…]
+Appendix C Ablation studies
+C.1 Ablation on Temperature
+We conduct an ablation study to analyze the effects of temperature on DISC and BoN.
+Temperature controls the randomness of token sampling in autoregressive models, influencing both exploration and consistency.
+Higher temperatures encourage more diverse outputs, whereas lower temperatures yield more deterministic generations.
+To examine its impact, we evaluate DISC and BoN on a 100-problem subset of APPS (the first 100 problems) using gpt-4o-mini.
+Figure˜16 presents the Pass@token scaling curve for DISC across different temperatures.
+The results indicate that lower temperatures lead to improved performance, as DISC benefits from more deterministic step selection.
+Unlike BoN, which relies on broad solution sampling, DISC dynamically refines steps, making stable token probabilities advantageous.
+Figure˜16 illustrates the frequency of actual partitions made by DISC at different temperatures.
+As temperature increases, the number of partitions fluctuates more, suggesting that high temperature introduces instability in step selection.
+Lower temperatures provide more structured decomposition, reducing unnecessary subdivisions.
+In Figure˜16, we visualize the mean reward per step.
+The trend shows a linear increase in reward as step number grows, demonstrating that deeper decomposition results in progressively better solutions.
+This reinforces that DISC effectively allocates computation towards refining difficult steps.
+The mean standard deviation per step is shown in Figure˜16.
+Lower temperatures yield lower standard deviations, confirming that DISC benefits from reduced variability in sample quality.
+This consistency allows for more reliable prioritization of difficult steps, enhancing overall inference efficiency.
+For comparison, Figure˜16 and Figure˜16 display Pass@token and Pass@k scaling curves for BoN across different temperatures.
+Unlike DISC, BoN achieves peak performance at a temperature around 0.6-0.8, balancing diversity and consistency.
+Higher temperatures increase exploration but degrade precision, while lower temperatures hinder sample diversity, reducing the probability of obtaining high-quality completions.
+These findings highlight the fundamental difference between DISC and BoN: DISC benefits from lower variance and stable decomposition, while BoN relies on broader exploration facilitated by moderate temperature settings.
+As a result, optimal temperature settings differ significantly between these methods, with DISC favoring deterministic sampling and BoN requiring a balance between diversity and coherence.
+Figure 11: Pass@token scaling curve for different temperatures on APPS using gpt-4o-mini. The lower the temperature, the stronger the DISC performance.
+Figure 12: Partition frequency of DISC with different temperatures on APPS using gpt-4o-mini.
+Figure 13: Mean reward per step of DISC with different temperatures on APPS using gpt-4o-mini. The mean reward scales linearly with step number.
+Figure 14: Mean standard deviation per step of DISC with different temperatures on APPS using gpt-4o-mini. Lower temperature means lower average standard deviation.
+Figure 15: Pass@k scaling curve for different temperatures on APPS using gpt-4o-mini for BoN. A temperature around 0.6–0.8 leads to the best performance and balance between diversity and consistency.
+Figure 16: Pass@token scaling curve for different temperatures on APPS using gpt-4o-mini for BoN.
+A temperature around 0.6–0.8 leads to the best performance and balance between diversity and consistency.
+C.2 Acceptance Method
+Figure 17: Token level comparison of different priority metrics on DISC in the APPS setting with gpt-4o-mini. Both Q and Z based priority metrics perform well.
+Figure 18: Partition frequency of DISC with different priority metrics on APPS using gpt-4o-mini.
+Figure 19: Mean reward per step of DISC with different priority metrics on APPS using gpt-4o-mini. All metrics display strong correlation between step depth and the mean reward.
+Figure 20: Mean standard deviation per step of DISC with different priority metrics on APPS using gpt-4o-mini. All metrics display correlation between step depth and the standard deviation.
+We analyze the impact of different acceptance methods on DISC’s performance.
+These methods govern whether a candidate prefix is accepted for further decomposition, directly influencing efficiency, stability, and final solution quality.
+We begin by introducing notation to characterize the distribution of rewards obtained when sampling completions from a given prefix $\boldsymbol{p}$.
+Let $R_{\boldsymbol{p}}$ denote the random variable of the reward of a completion sampled from the LLM policy $\pi$ conditioned on prefix $\boldsymbol{p}$:
+$R_{\boldsymbol{p}}:=R(\boldsymbol{p}\cdot\boldsymbol{s}),\quad\text{where }\boldsymbol{s}\sim\pi(\cdot\mid\boldsymbol{p}).$
+Let $F_{\boldsymbol{p}}$, $\mu_{\boldsymbol{p}}$, and $\sigma_{\boldsymbol{p}}$ be the cumulative distribution function (CDF), mean, and standard deviation of $R_{\boldsymbol{p}}$, respectively.
+To estimate these quantities in practice, we sample $M$ completions from the policy:
+$Y_{\boldsymbol{p}}=\left\{\left(\boldsymbol{p}\cdot\boldsymbol{s}^{i},\,R(\boldsymbol{s}^{i})\right)\;\middle|\;\boldsymbol{s}^{i}\sim\pi(\cdot\mid\boldsymbol{p})\right\}_{i=1}^{M}.$
+From these samples, we compute the empirical mean and standard deviation:
+$\mu_{\boldsymbol{p}}=\frac{1}{M}\sum_{i=1}^{M}R(\boldsymbol{s}^{i}),\quad\sigma_{\boldsymbol{p}}=\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left(R(\boldsymbol{s}^{i})-\mu_{\boldsymbol{p}}\right)^{2}}.$
+Let $r^{(1)}_{\boldsymbol{p}}=\max_{i}R(\boldsymbol{s}^{i})$ denote the sample maximum reward observed for prefix $\boldsymbol{p}$.
+We evaluate DISC on the first 200 competition-level APPS problems using gpt-4o-mini at a fixed temperature of 0.8. The following acceptance methods are compared:
+  • DISC-Z: accepts if the candidate prefix $\boldsymbol{c}$ has a lower z-score than the base prefix $\boldsymbol{b}$.
+  • DISC-Q: accepts if the candidate prefix has a lower mean reward, i.e., $\mu_{\boldsymbol{c}}<\mu_{\boldsymbol{b}}$.
+  • DISC-negZ: accepts if the candidate has a higher z-score.
+  • DISC-negQ: accepts if the candidate has a higher mean reward.
+  • DISC-R: accepts a candidate uniformly at random.
+Z-score Based Acceptance (DISC-Z).
+For a base prefix $\boldsymbol{b}$ and candidate prefix $\boldsymbol{c}$, we estimate the z-score of the sample maximum $r^{(1)}_{\boldsymbol{c}}$ as:
+$z_{\boldsymbol{c}}=\frac{r^{(1)}_{\boldsymbol{c}}-\mu_{\boldsymbol{c}}}{\sigma_{\boldsymbol{c}}},\quad\text{so }\mathbb{P}[R_{\boldsymbol{c}}>r^{(1)}_{\boldsymbol{c}}]=1-F_{\boldsymbol{c}}(r^{(1)}_{\boldsymbol{c}})$.
+We accept $\boldsymbol{c}$ if $z_{\boldsymbol{c}}<z_{\boldsymbol{b}}$. A lower z-score implies a greater tail probability mass and thus a higher chance of improvement.
+Q-based Acceptance (DISC-Q).
+This method accepts $\boldsymbol{c}$ if its sample mean is lower than that of $\boldsymbol{b}$:
+$\mu_{\boldsymbol{c}}<\mu_{\boldsymbol{b}}.$
+While simple, this method compares absolute expected values without accounting for reward variance, making it less robust to noise in $\pi$’s samples.
+Empirical Comparison.
+Figure˜20 shows token-level performance across methods. DISC-Z significantly outperforms alternatives, highlighting the value of normalizing reward advantage by variance.
+Figure˜20 shows that DISC-Z produces fewer but more meaningful partitions, suggesting better allocation of compute.
+Figure˜20 shows that DISC-Z achieves sharper reward gains early on, indicating better prioritization of impactful refinements.
+Figure˜20 reports step-wise reward variance. DISC-Z yields the most stable and consistent improvements.
+Why Z-score Works Best Z-score based acceptance balances the mean and variance of reward samples, estimating the probability of improvement in a statistically grounded way.
+This leads to more reliable and efficient decomposition decisions.
+Overall, our results confirm that acceptance method design is critical for dynamic decomposition. Among all tested methods, DISC-Z consistently delivers the best performance and compute efficiency.
+C.3 Model Ablation
+Figure 21: Pass@token scaling curve for open source models with DISC on APPS. DISC also demonstrates strong performance gains with open source models.
+Figure 22: Partition frequency of DISC with open source models on APPS.
+Figure 23: Mean reward per step of DISC with open source models on APPS.
+Figure 24: Mean standard deviation per step of DISC with open source models on APPS.
+We investigate how different LLMs perform when used with DISC on 200 competition-level APPS problems, given a sample budget of 30.
+The groundtruth reward model was used to evaluate correctness, and all models were set to a temperature of 0.8.
+Due to the challenging nature of the benchmark, open-source models struggled to achieve strong performance independently.
+However, when paired with DISC, their performance significantly improved.
+Figure˜24 presents the Pass@token scaling curve for open-source models using DISC.
+The results demonstrate that DISC substantially enhances the capabilities of these models, closing the gap between them and proprietary alternatives.
+Figure˜24 visualizes the partition frequency of DISC with different open-source models.
+Compared to their standalone performance, the use of DISC led to more structured and effective decomposition, highlighting its adaptability to different architectures.
+The mean reward per step is shown in Figure˜24.
+Similar to prior findings, we observe that deeper decomposition leads to increasingly higher rewards.
+Notably, even lower-capacity models benefit from DISC ’s ability to iteratively refine their solutions.
+Finally, Figure˜24 presents the mean standard deviation per step. With DISC, the variance in performance is significantly reduced, resulting in more stable and reliable inference.
+Overall, these findings emphasize that DISC is a robust framework capable of enhancing inference performance across diverse LLMs, particularly those with limited standalone capabilities.
+C.4 Ablation on Partition Fraction $\alpha$
+Figure 25: Token level comparison of different DISC splitting fraction $\alpha_{0}$ on APPS competition level. $0.15\leq\alpha_{0}\leq 0.25$ seems to be optimal.
+Figure 26: Effect of partition fraction $\alpha_{0}$ in DISC on APPS with gpt-4o-mini. The range $0.15\leq\alpha_{0}\leq 0.25$ appears optimal.
+We include additional analysis on the effect of the initial partition fraction $\alpha_{0}$, which determines the fraction of the best suffix used to propose a new candidate prefix at each iteration.
+As shown in Figures˜26 and 26, we observe that performance peaks in the range $0.15\leq\alpha_{0}\leq 0.25$, with both token-level and Pass@k metrics favoring this region.
+This behavior aligns with the underlying motivation of DISC: a smaller $\alpha_{0}$ results in more conservative candidate proposals—shorter steps that allow for finer-grained refinement.
+This is beneficial because committing to suboptimal prefixes early in the search can lead to irrevocable errors, especially in greedy search variants.
+Smaller values of $\alpha_{0}$ give the algorithm more flexibility to adjust course later on, while still making meaningful progress toward a complete solution.
+Conversely, large $\alpha_{0}$ values (e.g., $\alpha_{0}>0.3$) result in overly aggressive expansions that risk committing to noisy or premature completions, leading to reduced accuracy and inefficient use of sampling budget.
+Therefore, the optimal range of $\alpha_{0}$ reflects a balance between exploration and commitment—proposing candidate steps that are informative enough to guide search, yet cautious enough to preserve the ability to refine future decisions.
+This ablation confirms that adaptive decomposition benefits from conservative, incremental prefix extension when navigating complex reasoning or program synthesis tasks.
+C.5 Reward Distribution
+Figure 27: Reward distribution for 100 samples on a given MATH500 problem. The rewards appear to be roughly normal shaped.
+A key assumption in our algorithm is that the distribution of rewards for sampled completions from a given prefix follows a location-scale family, such as the Gaussian distribution.
+This assumption enables the use of z-scores to estimate the relative quality of candidate prefixes and to guide step acceptance.
+As shown in Figure˜27, the empirical reward distribution for 100 samples on a representative MATH500 problem appears approximately Gaussian, with a unimodal and symmetric shape.
+While we do not assume exact normality, this empirical observation supports the use of z-score-based comparisons, as the Gaussian approximation provides a reliable proxy for estimating tail probabilities and potential for reward improvement.
+We observe similar patterns across other problems and benchmarks, further validating this modeling choice.
+Appendix D Main Results Extended
+D.1 APPS
+Figure 28: Sampling frequency of each step averaged over the problems on APPS with gpt-4o-mini. DISC seems to have a slight preferences for spending more compute on earlier found steps.
+Figure 29: Comparison of actual and planned partitions on APPS.
+DISC outperforms other methods with fewer partitions by efficiently identifying critical steps.
+Unlike token and line split methods, which plan many partitions but search only a subset, DISC dynamically adjusts partitioning based on budget.
+We provide extended analysis of DISC ’s behavior on the APPS benchmark using gpt-4o-mini.
+Figure˜28 shows the average sampling frequency per step across all problems.
+Interestingly, DISC exhibits a slight preference for allocating more sampling budget to earlier-discovered steps.
+This behavior reflects the intuition that early steps in a solution often set up the structure for later reasoning, and refining these foundational steps yields more downstream improvements.
+Figure˜29 compares the number of actual partitions (i.e., accepted and committed steps) with the number of planned partitions (i.e., proposed candidate splits) across decomposition methods.
+Token-level and line-based methods typically predefine many steps but only explore a subset due to budget constraints, leading to inefficient allocation.
+In contrast, DISC adaptively proposes and accepts partitions in real time based on sample statistics and reward feedback, resulting in significantly fewer actual partitions while maintaining or exceeding performance.
+This efficiency demonstrates DISC ’s ability to identify and focus on high-impact steps, avoiding wasted compute on unnecessary refinements and enabling stronger performance within the same sampling budget.
+D.2 Additional Examples
+Below is another computed decomposition for the same problem as the one shown in the main text:
+DISC example decomposition (Lagrange method) [ Let $x$ be the length and $y$ be the width of the rectangle.
+Since the perimeter is 24 inches, we have $2x+2y=24\Rightarrow x+y=12$.
+] [ Therefore, ] [ we must maximize the area xyxy subject to this constraint.
+] [ We use the method of Lagrange multipliers.
+The Lagrangian is $\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}\mathcal{L}(x,y,\lambda)=xy-\lambda(x+y-12).$} We differentiate with respect to $x,y,$ and $\lambda$ to obtain $\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}\frac{\partial\mathcal{L}}{\partial x}=y-\lambda=0,\qquad\frac{\partial\mathcal{L}}{\partial y}=x-\lambda=0,\qquad\frac{\partial\mathcal{L}}{\partial\lambda}=x+y-12=0.$} We find that $x=y=\lambda$, so $x=y=\frac{12}{2}=6.$ Therefore, the maximum area of the rectangle is $6\cdot 6=\boxed{36}$.
+]
+DISC again highlights how initial constraint formulation is of medium importance—providing essential structure but not representing a reasoning fork.
+The token "Therefore," is marked as high-importance, capturing a critical conceptual transition from constraint setup to the optimization approach.
+Notably, the bulk of the mathematical machinery involving partial derivatives and substitutions receives low importance, consistent with the idea that such computations are largely mechanical once the decision to use Lagrange multipliers is made.
+Interestingly, the final step involving the boxed answer is also marked low, suggesting DISC allocates minimal compute here once earlier reasoning is settled—underscoring the idea that key inference pivots lie upstream in the decision flow.
+D.3 APPS with Self-generated Validation Tests
+We examine DISC performance on APPS when using self-generated validation tests.
+All methods utilized the same set of self-generated validation tests to ensure fair comparisons.
+Each problem received 5-10 validation tests, with the exact number determined dynamically by the LLM.
+We evaluated a subset of 100 APPS problems, generating samples until the sample budget was exhausted or a correct solution was found.
+Figure˜35 illustrates the Pass@token scaling curve, showing that DISC maintains strong scaling performance in this setting, though at a slightly lower rate compared to ground-truth verification.
+Figure˜35 and Figure˜35 compare actual and planned partition frequencies, respectively.
+The results indicate that DISC continues to make structured decompositions even with self-generated validation, preserving its efficiency.
+The mean reward per step, shown in Figure˜35, follows a similar trend as in previous experiments, reinforcing that DISC effectively allocates compute resources for iterative refinement.
+Lastly, Figure˜35 demonstrates that DISC maintains lower standard deviations in performance, indicating stable quality improvements across steps.
+Figure 30: Pass@k on APPS with gpt-4o-mini using self-generated validation tests.
+Figure 31: Token level comparison of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
+DISC still scales better than other methods in this setting, albeit at a lower rate.
+Figure 32: Actual partition frequency of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
+Figure 33: Planned partition frequency of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
+Figure 34: Mean reward per step of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
+Figure 35: Mean standard deviation different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
+D.4 MATH500
+Completions for MATH500 include both the reasoning steps and the final answer.
+Since MATH500 contains more problems than APPS200 and MATH problems tend to be relatively easier, solution quality saturates quickly.
+Therefore, we use a lower sample budget of 10 for these experiments.
+Figure˜39 presents the Pass@k performance for different decomposition methods on MATH500.
+We observe that all decomposition-based approaches achieve similar Pass@k performance, consistently outperforming BoN.
+This indicates that the structured nature of MATH problems allows multiple decomposition strategies to be effective.
+Despite similar Pass@k results, the true advantage of DISC lies in its token efficiency, as shown in Figure˜5.
+DISC significantly reduces the number of tokens required to reach correct solutions compared to alternative methods, demonstrating its ability to allocate computational effort efficiently in mathematical reasoning tasks.
+Additionally, we analyze the partitioning behavior of DISC on MATH500.
+Figure˜39 illustrates the actual partition frequency for different decomposition methods.
+The planned partitioning behavior, shown in Figure˜39, further highlights how DISC effectively balances exploration and refinement.
+Finally, we present the mean standard deviation per step in Figure˜39.
+Lower variance suggests that DISC produces more stable and reliable decompositions over multiple runs, reinforcing its robustness in both mathematical and program synthesis domains.
+Figure 36: Pass@k performance comparison for different decomposition methods on MATH500. DISC consistently outperforms BoN across different sampling budgets.
+Figure 37: Observed partition frequency of different decomposition methods on MATH500. DISC effectively segments problems into meaningful subcomponents.
+Figure 38: Planned partitioning strategy of different decomposition methods on MATH500. DISC ’s structured approach leads to more efficient problem breakdowns.
+Figure 39: Mean standard deviation per step for different decomposition methods on MATH500. Lower variance in DISC suggests more stable and reliable problem-solving steps.
+D.5 LiveCodeBench
+We evaluate DISC on LiveCodeBench, a benchmark designed for code generation tasks with a focus on real-world software development challenges.
+LiveCodeBench presents a unique set of problems requiring both reasoning and structured decomposition, making it a suitable testbed for evaluating DISC’s ability to refine and improve intermediate steps.
+Figure˜43 shows the Pass@k comparison of different decomposition methods on LiveCodeBench.
+DISC consistently scales better than other decomposition methods, highlighting its ability to refine intermediate steps more effectively in complex coding scenarios.
+Figure˜43 illustrates the observed partition frequency of different decomposition methods.
+The structured approach of DISC results in well-balanced decomposition across steps, reducing unnecessary partitioning while maintaining sufficient granularity for improved solution refinement.
+Figure˜43 displays the planned partition frequency across methods.
+DISC dynamically determines the most effective partitions based on the evolving problem state, leading to more targeted and efficient decompositions.
+Finally, Figure˜43 presents the mean standard deviation per step across decomposition methods.
+Lower variance in DISC suggests that it produces more stable and reliable decompositions, reinforcing its robustness for solving LiveCodeBench problems.
+Figure 40: Pass@k performance comparison for different decomposition methods on LiveCodeBench. DISC consistently outperforms other methods in structured problem refinement.
+Figure 41: Observed partition frequency of different decomposition methods on LiveCodeBench. DISC effectively balances problem segmentation while avoiding excessive partitioning.
+Figure 42: Planned partitioning strategy of different decomposition methods on LiveCodeBench. DISC dynamically adapts its partitioning to optimize search efficiency.
+Figure 43: Mean standard deviation per step for different decomposition methods on LiveCodeBench. Lower variance in DISC suggests more stable and reliable problem-solving steps.
+D.6 Computational Overhead
+We analyze the computational overhead of DISC by measuring the runtime breakdown between actual LLM token generation and all auxiliary operations (e.g., z-score computation, candidate prefix management, and recursive decomposition logic).
+Since token generation is the dominant cost in LLM inference, reporting relative percentages of runtime—rather than absolute wall-clock time—offers a more meaningful and consistent comparison.
+Absolute runtime can vary significantly across hardware configurations, backend optimizations, or batching strategies, whereas percentage-based measurements isolate algorithmic overhead from infrastructure-dependent variance.
+As shown in Figure˜7, the proportion of runtime DISC spends on LLM token generation versus auxiliary logic closely matches that of baseline methods such as BoN and LineSplit.
+This indicates that DISC introduces negligible additional overhead despite its dynamic behavior and internal scoring computations.
+In practice, more than 90% of total runtime is still dominated by LLM token sampling across all methods.
+Furthermore, because DISC achieves better performance with fewer tokens (i.e., improved token efficiency), its effective runtime per successful solution is actually lower than that of less efficient baselines.
+This combination—minimal added overhead and superior token scaling—makes DISC a highly practical method for real-world inference settings, where total compute cost is often tightly constrained.
+Appendix E Limitations
+While DISC demonstrates strong empirical performance and generality across tasks and models, several limitations merit discussion.
+Reliance on LLM-Generated Test Validation.
+Our evaluation framework employs self-generated test validation, where the LLM is prompted to produce its own unit tests for a given code generation task, and these tests are used as a proxy reward model.
+This approach enables scalable evaluation in the absence of ground-truth test cases, but its effectiveness depends critically on the quality of the generated tests.
+In particular:
+  • The LLM’s ability to produce meaningful and comprehensive tests directly impacts evaluation fidelity. Poor coverage or semantically shallow tests may overestimate correctness.
+  • The framework assumes that most generated tests are correct and executable. For complex or underspecified tasks, this assumption may fail, leading to false positives or noisy reward signals.
+  • Since generated tests may not align with reference solutions, cross-method comparison can become inconsistent.
+These issues are consistent with observations in prior work such as CodeT [30], which highlights tradeoffs between scalability and reliability in test-based self-evaluation.
+Dependence on Reward Model Availability.
+DISC requires access to a scalar reward signal to guide step-wise decomposition.
+While tasks like code generation or math reasoning naturally provide verifiers (e.g., test cases or numerical checks), applying DISC to domains lacking explicit reward signals necessitates constructing auxiliary reward models or LLM-based critics, which introduces additional complexity and potential bias.
+Scope Limited to Single-Turn Generation.
+The current formulation of DISC is designed for single-pass generation tasks.
+It does not yet handle multi-turn or interactive settings where intermediate reasoning steps can elicit feedback or modify the problem context.
+Extending DISC to such interactive regimes remains an important direction for future work.
+Reduced Benefit in Trivial or Non-Compositional Tasks.
+DISC allocates computational budget adaptively across reasoning steps.
+When a task is trivially solvable or lacks compositional structure—such that early reasoning does not meaningfully constrain the final outcome—the gains from decomposition and step-wise control diminish.
+Despite these limitations, we believe that DISC provides a strong foundation for adaptive, reward-driven reasoning.
+Future extensions could integrate learned verifiers, support multi-turn interaction, and explore joint reasoning–evaluation co-evolution.
+Appendix F Search and Scaling
+F.1 DISC Plug-and-Play Search
+Algorithm 3 DISC: Decomposition for Plug-and-Play Search
+1:LLM $\pi$, Reward model $R$, prompt $x$, initial partition fraction $\alpha_{0}$, negative binomial threshold $\sigma$, total budget $N$, current budget $n$
+2:
+3:function ExpandNode(parent prefix $\boldsymbol{p}_{b}$, parent z score $z_{b}$, suffix to child $\boldsymbol{s}_{c}$)
+[…]
+5:while $n<N$ do
+6:  $\boldsymbol{p}_{c}=\boldsymbol{p}_{b}\cdot\textup{split}(\boldsymbol{s}_{c},\alpha)$
+7:  $Y_{c}=$MakeChildren($\boldsymbol{p}_{c}$)
+8:  $z_{c}=(\text{max}(Y_{c})-\text{mean}(Y_{c}))/\text{std}(Y_{c})$
+9:  $n\leftarrow n+M$
+10:  if $z_{c}<z_{b}$ then
+11:   break
+12:  else
+13:   $\alpha\leftarrow\alpha_{0}\alpha$
+14:  end if
+15:end while
+16:return $\boldsymbol{p}_{c}$, $Y_{c}$
+17:
+18:function MakeChildren(parent prefix $\boldsymbol{p}_{b}$)
+[…]
+20:return Y
+F.2 Monte Carlo Tree Search (MCTS)
+Monte Carlo Tree Search (MCTS) is a widely used algorithm for sequential decision-making in large search spaces, particularly in applications such as game playing, planning, and inference scaling.
+The algorithm builds a search tree incrementally by simulating different sequences of actions and updating estimates of state quality.
+A key advantage of MCTS is its ability to balance exploration (discovering new states) and exploitation (refining promising ones) using a data-driven search process.
+The MCTS pipeline consists of four fundamental steps: selection, expansion, simulation, and backpropagation.
+F.2.1 Selection
+Starting from the root node representing the current state $\boldsymbol{s}$, MCTS iteratively traverses the search tree by selecting child nodes based on a selection policy.
+The most commonly used selection criterion is the Upper Confidence Bound for Trees (UCT), which balances exploration and exploitation:
+$UCT(\boldsymbol{s},\boldsymbol{d})=\hat{Q}(\boldsymbol{s},\boldsymbol{d})+c\sqrt{\frac{\ln\left(\sum_{\boldsymbol{b}}n(\boldsymbol{s},\boldsymbol{b})\right)}{n(\boldsymbol{s},\boldsymbol{d})}},$ (1)
+where $\hat{Q}(\boldsymbol{s},\boldsymbol{d})$ represents the estimated value of selecting action $\boldsymbol{d}$ from state $\boldsymbol{s}$, $n(\boldsymbol{s},\boldsymbol{d})$ is the visit count for this action, and $c$ is a hyperparameter controlling the trade-off between exploring new actions and favoring those with high past rewards.
+F.2.2 Expansion
+Once a leaf node (a previously unexplored state) is reached, the algorithm expands the tree by adding one or more new nodes.
+These new nodes represent potential future states $\boldsymbol{s}^{\prime}$ generated by sampling an action $\boldsymbol{d}$ from a predefined policy.
+This step broadens the search space and allows MCTS to evaluate new possibilities.
+F.2.3 Simulation
+Following expansion, the algorithm conducts a simulation (or rollout) from the newly added state.
+This step involves generating a sequence of actions according to a predefined policy until reaching a terminal state or an evaluation horizon.
+The outcome of the simulation, denoted as $v(\boldsymbol{s}^{\prime})$, provides an estimate of the quality of the new state.
+Depending on the application, this could represent a game result, an optimization score, or an inference accuracy metric.
+F.2.4 Backpropagation
+The final step involves propagating the results of the simulation back up the search tree to refine the estimated values of prior states and actions.
+Each node along the trajectory $\tau=[\boldsymbol{s}_{0},\boldsymbol{d}_{1},\boldsymbol{s}_{2},\dots,\boldsymbol{s}_{-1}]$ is updated iteratively:
+$\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t+1)}\leftarrow(1-\alpha_{n})\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t)}+\alpha_{n}\max\{\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t)},\hat{Q}(\boldsymbol{s}_{i+1},\boldsymbol{d}_{i+2})^{(t+1)}\},$ (2)
+where $\alpha_{n}$ is a learning rate that depends on the visit count, and the maximum function ensures that the best-performing trajectories are emphasized.
+MCTS has been widely adopted in inference scaling techniques due to its ability to efficiently allocate computational resources, focusing more on high-reward states while avoiding unnecessary exploration of unpromising regions.
+In later sections, we explore how MCTS can be combined with dynamic decomposition to further optimize inference scaling.
+F.2.5 Combining Dynamic Decomposition with MCTS
+MCTS can be enhanced by integrating dynamic decomposition, where each node in the search tree represents a decomposition of the problem into steps.
+Instead of treating states as atomic decisions, we recursively decompose reasoning steps, dynamically adjusting granularity based on difficulty.
+In this framework:
+  • Each node in the MCTS tree represents a partial decomposition of the problem, with child nodes corresponding to alternative step partitions.
+  • Branching occurs by generating candidate next steps using dynamic decomposition, allowing finer steps for complex regions while maintaining efficiency for simpler ones.
+  • The selection step prioritizes nodes that represent more promising decompositions, dynamically refining challenging areas through recursive subdivision.
+  • The backpropagation step ensures that decompositions leading to high-quality solutions are reinforced, helping the search tree converge toward optimal inference paths.
+By integrating dynamic decomposition with MCTS, we efficiently allocate compute to the most critical reasoning steps, improving inference quality while maintaining computational efficiency.
+F.3 Beam Search
+Beam search is a heuristic search algorithm commonly used in inference tasks where computational efficiency is a priority.
+Unlike exhaustive search methods, beam search maintains only the top $k$ best candidates at each step, making it an effective strategy for structured prediction problems and sequential decision-making.
+At each iteration:
+  • The algorithm selects the $k$ most promising partitions from the previous step based on an evaluation metric.
+  • Each selected partition is expanded by generating possible next-step samples.
+  • The newly generated partitions are ranked, and only the top $k$ candidates are retained for the next iteration.
+  • This process continues until a stopping criterion is met, such as reaching a predefined depth or finding a sufficiently high-quality solution.
+Beam search provides a computationally efficient way to explore structured solution spaces while maintaining high-quality search trajectories.
+By integrating beam search with dynamic decomposition, we ensure that inference computation is allocated efficiently, focusing on the most promising reasoning paths at each step.
+F.4 Additional Results and Analysis
+Figure 44: Pass@k on APPS with gpt-4o-mini using different search methods. MCTS scales best, followed by greedy search, followed by beam search.
+Figure 45: Token level comparison of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. MCTS scales best, followed by greedy search, followed by beam search.
+Figure 46: Actual partition frequency of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. Greedy is able to search to higher depths given the same sampling budget.
+Figure 47: Mean standard deviation of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. All search methods display decreasing standard deviation with search depth.
+Experiments comparing different search methods were conducted on a 100-problem subset of the APPS dataset (first 100 problems) using GPT-4o-mini.
+All methods used a temperature of 0.2, with $\alpha=0.15$, Q priority metric, and $\sigma=1.0$.
+Token-level comparison: As shown in Figure 47, MCTS scales best among the tested methods, demonstrating superior efficiency in identifying promising partitions.
+Greedy search follows closely, while beam search exhibits the slowest scaling.
+Partition frequency analysis: Figure 47 reveals that greedy search explores to greater depths within the same sampling budget.
+This suggests that greedy search prioritizes deep refinements, whereas MCTS and beam search balance depth with breadth.
+Step variance analysis: Figure 47 illustrates that all search methods display decreasing standard deviation with increasing search depth.
+This trend indicates that deeper searches converge towards stable, high-quality partitions, reinforcing the benefits of dynamic decomposition.
+These results highlight the trade-offs between search methods: MCTS offers robust exploration-exploitation balance, greedy search favors depth-first refinement, and beam search provides a structured yet computationally constrained approach.
+The integration of dynamic decomposition further enhances these search strategies by adaptively allocating computational resources to critical reasoning steps.
+Appendix G Theoretical analysis
+G.1 Main Theorem
+We begin by introducing notation to characterize the distribution of rewards obtained when sampling completions from a given prefix $\boldsymbol{p}$.
+Let $R_{\boldsymbol{p}}$ denote the random variable of the reward of a completion sampled from the LLM policy $\pi$ conditioned on prefix $\boldsymbol{p}$:
+$R_{\boldsymbol{p}}:=R(\boldsymbol{p}\cdot\boldsymbol{s}),\quad\text{where }\boldsymbol{s}\sim\pi(\cdot\mid\boldsymbol{p}).$
+Let $F_{\boldsymbol{p}}$, $\mu_{\boldsymbol{p}}$, and $\sigma_{\boldsymbol{p}}$ be the cumulative distribution function (CDF), mean, and standard deviation of $R_{\boldsymbol{p}}$, respectively.
+We define a solution $\boldsymbol{y}^{*}$ as correct if it achieves a reward of at least $r^{*}=1$, i.e., $R(\boldsymbol{y}^{*})\geq 1$.
+Then, the probability that a random sample from $\pi(\cdot\mid\boldsymbol{p})$ yields a correct solution is given by:
+$\mathbb{P}[R_{\boldsymbol{p}}\geq 1]=1-F_{\boldsymbol{p}}(1).$
+For a given set of $M$ independent samples from $\pi(\cdot\mid\boldsymbol{p})$, let $r^{\max}_{\boldsymbol{p}}$ denote the maximum observed reward among them, i.e., the sample maximum.
+We make the following problem assumption:
+Assumption 1
+A correct solution exists in the support of the policy $\pi(\boldsymbol{y}^{*}|\boldsymbol{x})>0$ and the length of the correct solution is finite: $|\boldsymbol{y}^{*}|=K<\infty$.
+We make the following assumptions about the policy $\pi$:
+Assumption 2
+We assume $R_{\boldsymbol{p}}$ belongs to a location-scale family with base CDF $F_{R}$.
+Remark 2
+The location-scale family includes common distributions such as the Normal, Cauchy, and Logistic distributions, which are characterized by being affine transformations of a fixed base distribution.
+Specifically, a random variable $X$ belongs to a location-scale family if it can be written as $X=\mu+\sigma Z$, where $Z\sim F_{R}$ is a standardized random variable, and $\mu\in\mathbb{R}$ and $\sigma>0$ are the location and scale parameters, respectively.
+We observe empirical evidence supporting this modeling choice for $R_{\boldsymbol{p}}$ in Section˜C.5.
+Assumption 3 (Reward distribution converges with prefix)
+Consider some base prefix $\boldsymbol{b}$ and some candidate prefix $\boldsymbol{c}=\boldsymbol{b}\cdot\boldsymbol{x}$ for any token sequence $\boldsymbol{x}$.
+Then, $\sigma_{\boldsymbol{c}}\leq\sigma_{\boldsymbol{b}}$.
+Furthermore, if $\boldsymbol{b}$ is a complete solution (it ends in an EOS token) then $\sigma_{\boldsymbol{b}}=0$.
+Remark 3
+This assumption reflects the intuition that as more tokens are appended to a prefix, the LLM becomes increasingly committed to a narrower set of likely continuations, leading to lower uncertainty in the reward distribution.
+In the limit, once a complete solution is formed (i.e., the sequence ends with an EOS token), the reward becomes deterministic, and the variance collapses to zero.
+We emphasize that this assumption is not required for proving the optimality of DISC.
+However, the greater the difference between $\sigma_{\boldsymbol{c}}<\sigma_{\boldsymbol{b}}$, the faster the convergence, as DISC will be able to prune suboptimal paths more confidently with fewer samples.
+Assumption 4 (Reward distribution converges slowly)
+If Algorithm˜1 accepts a candidate prefix $\boldsymbol{c}$ over a base prefix $\boldsymbol{b}$, then we assume the standard deviation of the reward distribution under $\boldsymbol{c}$ is a lower bounded:
+$\sigma_{\boldsymbol{c}}\geq\left(1-\frac{\delta\sigma_{\boldsymbol{c}}}{\Delta}\right)\sigma_{\boldsymbol{b}}$
+where $\Delta=r^{*}-r^{\max}_{\boldsymbol{c}}$ is the optimality gap between the true maximum reward $r^{*}$ and the current best sample under prefix $\boldsymbol{b}$, and $\delta=\frac{r^{\max}_{\boldsymbol{b}}-\mu_{\boldsymbol{b}}}{\sigma_{\boldsymbol{b}}}-\frac{r^{\max}_{\boldsymbol{c}}-\mu_{\boldsymbol{c}}}{\sigma_{\boldsymbol{c}}}$ is the z-score drop, which quantifies the change in how surprising the best sample is under each distribution.
+Remark 4
+This assumption enforces that if we transition from base prefix $\boldsymbol{b}$ to candidate prefix $\boldsymbol{c}$, the standard deviation (spread) of the reward distribution under $\boldsymbol{c}$ cannot shrink too drastically—especially when: (i) we are still far from the optimal reward (i.e., $\Delta$ is large), and (ii) the improvement in sample quality is not significant (i.e., $\delta$ is small).
+Intuitively, this prevents committing to a sharp but unreliable prefix unless there is strong evidence of progress.
+If $\boldsymbol{c}$ does not yield sufficiently better samples or has low variance, it may hinder further exploration.
+This assumption ensures that accepted prefixes preserve enough reward diversity to support continued search.
+Remark 5
+While we list the confidence property ˜4 as an assumption, it can also be enforced directly as an explicit acceptance criterion in Algorithm˜1.
+Specifically, the algorithm could be modified to accept a candidate prefix $\boldsymbol{c}$ over base $\boldsymbol{b}$ only if both $\delta\geq 0$ and: $\left(1-\frac{\delta\sigma_{\boldsymbol{c}}}{\Delta}\right)\sigma_{\boldsymbol{b}}\leq\sigma_{\boldsymbol{c}}.$ where $\Delta$ can be computed because $r^{*}=1$.
+Combining Assumptions 3 and 4, we can write: $\frac{\sigma_{c}}{\sigma_{b}}\in\left[\left(1-\frac{\delta\sigma_{\boldsymbol{c}}}{\Delta}\right),1\right]$. This balances exploration with exploitation.
+Assumption 5 (Accurate estimates of $\mu_{\boldsymbol{p}}$ and $\sigma_{\boldsymbol{p}}$ )
+We assume that we sample $R_{\boldsymbol{p}}$ enough times to get accurate estimates of the true mean $\hat{\mu}_{\boldsymbol{p}}\approx\mu_{\boldsymbol{p}}$ and standard deviation $\hat{\sigma}_{\boldsymbol{p}}\approx\sigma_{\boldsymbol{p}}$.
+Because estimates of mean and standard deviation are accurate, the estimates of the z-scores are also accurate.
+Next we prove that Alg. 1 will never commit a candidate prefix that with a lower probability of sampling a correct solution than the base.
+Lemma 1
+Consider Algorithm 1, applied to a base prefix $\boldsymbol{b}$ and a candidate prefix $\boldsymbol{c}$, with corresponding reward distributions $F_{\boldsymbol{b}}$ and $F_{\boldsymbol{c}}$.
+Then, the probability of sampling a correct suffix from $\pi$ does not decrease after accepting $\boldsymbol{c}$:
+$\pi(\boldsymbol{s}^{*}\mid\boldsymbol{b})\leq\pi(\boldsymbol{s}^{*}\mid\boldsymbol{c}).$
+Proof 1
+A candidate prefix $\boldsymbol{c}$ is accepted by Algorithm 1 if either of the two conditions are true: (i) the standardized score (z-score) of $\boldsymbol{c}$ is lower than that of $\boldsymbol{b}$: $z_{c}<z_{b}$ or (ii) the first $\alpha$ fraction of tokens in the best suffix contains no tokens, and the prefix is accepted by default.
+In case (ii): if $\boldsymbol{c}$ is the empty prefix, then $\boldsymbol{c}=\boldsymbol{b}$, and the distributions are unchanged, so the result follows trivially.
+It remains to show case (i): if $z_{c}<z_{b}$ then, $\pi(\boldsymbol{s}^{*}\mid\boldsymbol{c})\geq\pi(\boldsymbol{s}^{*}\mid\boldsymbol{b})$.
+First, we define the z-score change $\delta>0$ as $\delta=\frac{r^{\max}_{\boldsymbol{b}}-\mu_{\boldsymbol{b}}}{\sigma_{\boldsymbol{b}}}-\frac{r^{\max}_{\boldsymbol{c}}-\mu_{\boldsymbol{c}}}{\sigma_{\boldsymbol{c}}}$ and the optimality gap $\Delta>0$ as $\Delta=r^{*}-r^{\max}_{\boldsymbol{c}}$.
+Then, we note that our algorithmic implementation initializes the set of candidate samples with the best solution, i.e.
+$\text{max}(Y_{b})\in Y_{c}$.
+Therefore, we have that the sample max of the candidate rewards is greater than or equal to that of the base: $r_{\boldsymbol{c}}^{\max}\geq r_{\boldsymbol{b}}^{\max}$.
+Using this information, we can rewrite the z-score inequality:
+$\frac{r^{*}-\mu_{\boldsymbol{b}}}{\sigma_{\boldsymbol{b}}}-\frac{r^{*}-\mu_{\boldsymbol{c}}}{\sigma_{\boldsymbol{c}}}$ $\geq\left(\delta-\frac{\Delta}{\sigma_{\boldsymbol{c}}}+\frac{\Delta}{\sigma_{\boldsymbol{b}}}\right)\geq 0$ (3)
+where we first plug in the definitions of $\delta$ and $\Delta$, do algebraic manipulation, and then apply assumption 4.
+The final inequality implies $F_{\boldsymbol{c}}(r^{*})<F_{\boldsymbol{b}}(r^{*})$, which implies $\pi(\boldsymbol{s}^{*}\mid\boldsymbol{c})\geq\pi(\boldsymbol{s}^{*}\mid\boldsymbol{b})$ because $\mathbb{P}[R_{\boldsymbol{p}}\geq r^{*}]=1-F_{\boldsymbol{p}}(r^{*})$ and $F_{\boldsymbol{p}}$ is monotonic in z score.
+Theorem 2 (Optimality of DISC)
+Consider Algorithm 1 applied to a problem input $\boldsymbol{x}$ and assume Assumptions 1, 2, 4, and 5 hold.
+Then, with probability 1, there exists a finite number of accepted prefixes $k>0$ and algorithm iterations $n>0$ such that the algorithm terminates, and the best solution $\boldsymbol{y}_{k}$ found is a correct solution: $R(\boldsymbol{y}_{k})\geq 1$.
+Proof 2
+We define the base prefix $\boldsymbol{b}_{k}$ as the result of appending $k$ accepted candidate prefixes. We now proceed with the induction.
+We first prove by induction that the probability of sampling a correct suffix remains strictly positive after each accepted prefix.
+Base Case ($k=0$): Initially, $\boldsymbol{b}_{0}=\boldsymbol{x}$.
+By assumption, there exists at least one correct solution $\boldsymbol{y}^{*}=\boldsymbol{x}\cdot\boldsymbol{s}^{*}$ such that $\pi(\boldsymbol{s}^{*}\mid\boldsymbol{x})>0$ and $R(\boldsymbol{y}^{*})\geq 1$.
+Hence, there is a non-zero probability of sampling a correct suffix from $\boldsymbol{b}_{0}$.
+Inductive Step: Assume that after $k$ accepted prefixes, the base prefix $\boldsymbol{b}_{k}$ satisfies:
+$\pi(\boldsymbol{s}^{*}\mid\boldsymbol{b}_{k})>0\quad\text{and}\quad R(\boldsymbol{b}_{k}\cdot\boldsymbol{s}^{*})\geq 1$
+for some correct suffix $\boldsymbol{s}^{*}$.
+Let $\boldsymbol{b}_{k+1}=\boldsymbol{b}_{k}\cdot\boldsymbol{c}$ be the base prefix after appending the $(k+1)$-th accepted candidate prefix $\boldsymbol{c}$.
+By Lemma 1, the probability of sampling a correct suffix does not decrease:
+$\pi(\boldsymbol{s}^{*}\mid\boldsymbol{b}_{k+1})\geq\pi(\boldsymbol{s}^{*}\mid\boldsymbol{b}_{k})>0.$
+Therefore, at every accepted prefix $\boldsymbol{b}_{k}$, the probability of sampling a correct suffix remains bounded below by some $\varepsilon>0$, where $\varepsilon\geq\pi(\boldsymbol{y}^{*}\mid\boldsymbol{x})$.
+Termination: Let $M_{k}$ be the number of samples from accepted candidate prefix $\boldsymbol{b}_{k}$.
+Since we sample from the accepted candidate prefix $\boldsymbol{b}_{k}$ at least once ($M_{k}\geq 1$), samples are independent, and the probability of sampling a correct suffix is at least $\pi(\boldsymbol{y}^{*}\mid\boldsymbol{x})>0$, the probability of not sampling a correct solution after accepting $k$ prefixes is at most
+$\left(1-\pi(\boldsymbol{y}^{*}\mid\boldsymbol{x})\right)^{k}.$
+Hence, the probability of never sampling a correct solution after infinitely many accepted prefixes is
+$\lim_{k\to\infty}\left(1-\pi(\boldsymbol{y}^{*}\mid\boldsymbol{x})\right)^{k}=0,$
+which implies that, with probability 1, the algorithm will eventually sample a correct suffix.
+Furthermore, this implies that the number of accepted prefixes $k$ before obtaining a correct solution is almost surely finite.
+This follows from the fact that the number of independent trials before the first success with fixed success probability $p:=\pi(\boldsymbol{y}^{*}\mid\boldsymbol{x})>0$ is a geometric random variable, which is finite with probability 1.
+Therefore, there exists a finite number of accepted prefixes $k>0$ such that the algorithm terminates and returns a correct solution $\boldsymbol{y}^{*}$ satisfying $R(\boldsymbol{y}^{*})\geq 1$.
+Since the algorithm either accepts a candidate prefix or contracts it by a factor $\alpha\in(0,1)$ until the prefix length is 0, there can only be finitely many contractions before the candidate prefix is rejected or accepted.
+Therefore the algorithm must accept a candidate prefix in finite number of iterations, and so the total number of algorithm iterations $n$ before finding a correct solution is also finite.
+Remark 6
+Unlike Best-of-N (BoN), where the probability of sampling a correct solution remains constant across samples, DISC dynamically refines the sampling distribution by appending informative prefixes.
+As established in Lemma 1, the probability of generating a correct solution under DISC is non-decreasing across iterations.
+Consequently, DISC achieves faster convergence to a correct solution in expectation compared to BoN.
+This theoretical advantage is corroborated by our empirical results in Figure˜5, which show that DISC consistently outperforms BoN in terms of efficiency and scalability with respect to token budget.
+G.2 A Motivating Example on DISC
 We use the Wiener process $W(t)$ as an example where there are intractably many actions and steps.
@@
 Alternatively, suppose there is a promising path $w(\cdot)$ with a high final reward $w(T)=R$.
-It is natural to consider starting at some midpoint $\alpha T$ ($10<\alpha<$ 1) and perform more completions to obtain an even higher reward than $R$.
+It is natural to consider starting at some midpoint $\alpha T$ ($0<\alpha<1$) and perform more completions to obtain an even higher reward than $R$.
 The reward distribution sampled this way is
@@
 $\mathbb{P}(W^{\prime}(T)>R)=1-\Phi\bigg(\frac{R-w(\alpha T)}{\sqrt{(1-\alpha)T}}\bigg).$
-3.5 Example Decomposition
-With sample budget 100, the decomposition of a MATH problem is as follows, where color indicates the value of the priority metric $h$ of each step (yellow low, purple high).
-DISC example decomposition Let the length of the rectangle be $l$ and the width of the rectangle be $w$.
-Since the perimeter of the rectangle is 24 inches, we have that $242l+2w=$ 24, so $l+w=12$.
-We wish to maximize the area of the rectangle, which is $A=lw$.
-Let $l=12-w$ and plug into the area: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}A=(12-w)w\Rightarrow\qquad A=12w-w^{2}}$ Now, we differentiate $A$ with respect to $w$: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}A^{\prime}(w)=12-2w}$ We wish to maximize $A$, so we set $A^{\prime}(w)=0$, and solve for $w$: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}12-2w=0\Rightarrow\qquad w=6}$ Since $l=12-w$, we have that $l=12-6=6$.
-Therefore, the area of the rectangle is $A=lw=6\cdot 6=\boxed{36}$.
-Once the LLM generates the first three steps, the rest is easy. Interestingly, ‘which’ is an important decision point which helps decide how the LLM will complete the solution.
-Takeaway: Autoregressive models require autoregressive decomposition While words such as ‘which’, ‘therefore’, etc.
-may not seem like important steps to humans, they actually represent important steps for autoregressive LLMs which are trained on next token prediction.
-4 Experimental Results
-4.1 Benchmarks
-We evaluate DISC on three benchmarks: APPS, MATH, and LiveCodeBench, to assess its impact on inference scaling for both coding and reasoning.
-APPS (Hendrycks et al., 2021a) consists of 5000 competitive programming problems across three difficulty levels, with the competition-level subset being the hardest.
-We evaluate on a 200-problem subset due to computational constraints.
-MATH (Hendrycks et al., 2021b) comprises 12,500 math problems.
-Since the ground-truth verifier provides only binary rewards, we use a pretrained ORM (Xiong et al., 2024), trained via the method in (Wang et al., 2024b), with Llama-3.1-8B-Instruct as the base model.
-We test on a 500-problem subset (MATH500), identical to prior work (Wang et al., 2024b; Lightman et al., 2023).
-LiveCodeBench (Jain et al., 2024) is a continuously updated dataset from Leetcode, AtCoder, and CodeForces, ensuring LLMs have not been exposed to test problems.
-We evaluate on the 108 problems uploaded between 10/01/2024 and 12/01/2024 to prevent contamination.
-Figure 5: Token-level comparison on APPS competition level. DISC consistently outperforms prior decomposition methods in efficiency.
-Figure 6: Token-level comparison on MATH500. DISC outperforms other decomposition methods, demonstrating its effectiveness in mathematical reasoning.
-Figure 7: Token-level comparison on LiveCodeBench. DISC achieves superior inference scaling over TokenSplit, LineSplit, and BoN.
-4.2 Decomposition Comparison
-We compare DISC against three prior decomposition methods from Sec.
-2.2: TokenSplit (token-level decomposition), LineSplit (newline-based decomposition), and BoN (treating the entire solution as a single step).
-Across all benchmarks, DISC achieves superior scaling and performance under both fixed token budgets (Fig.
-7) and sample budgets (Fig.
-12).
-We evaluate two key metrics: Pass@k, the proportion of problems solved within a sample budget $k$, and Pass@token, the proportion solved within a given token budget.
-Notably, DISC consistently outperforms static decomposition methods on APPS, MATH, and LiveCodeBench (Fig.
-567), demonstrating its ability to allocate compute adaptively for improved inference efficiency.
-Extended results and analyses for each benchmark are provided in App.
-D.1D.4, and D.5.
-4.3 Decomposition Analysis and Interpretation
-Our results strongly indicate that decomposition—whether line-based, token-based, or DISC —improves sample quality.
-Figures 8 and 9 illustrate how the mean and variance of sampled rewards evolve with the step number, which represents the order in which a step is explored.
-Higher step numbers correspond to deeper search levels, where solutions are partitioned into finer-grained steps.
-As shown in Fig.
-8, increasing step number correlates with higher-quality solutions, demonstrating that finer-grained decomposition improves sample quality.
-Additionally, Fig.
-9 shows that reward variance decreases as step count increases, highlighting how decomposition enhances sampling precision.
-Furthermore, DISC achieves better performance with fewer partitions under a fixed sampling budget (Fig.
-10).
-We distinguish between actual partitions, the number of steps effectively explored, and planned partitions, the number of partitions intended by the method.
-Token and line split methods generate a large number of planned partitions (Fig.
-10) but search over at most 15 steps due to budget constraints.
-In contrast, DISC dynamically adjusts the number of partitions based on available budget, efficiently identifying and focusing on critical steps.
-Takeaway: Decomposition and sample quality Finer-grained decomposition improves sampled solution quality and reduces reward variance.
-Figure 8: Average reward per step, averaged over APPS problems.
-From step 3 onward, higher step counts strongly correlate with increased average reward, demonstrating the effectiveness of decomposition.
-The dip between steps 1 and 3 likely occurs because simple problems are solved early, preventing further search.
-Figure 9: Standard deviation of rewards per step, averaged over APPS problems. Decomposition reduces sampling variance, improving precision at deeper search depths.
-Figure 10: Comparison of actual and planned partitions on APPS.
-DISC outperforms other methods with fewer partitions by efficiently identifying critical steps.
-Unlike token and line split methods, which plan many partitions but search only a subset, DISC dynamically adjusts partitioning based on budget.
-4.4 Interaction Between Temperature and DISC
-Figure 11: Effect of temperature on DISC performance on APPS with gpt-4o-mini. Unlike BoN and other inference scaling methods, DISC achieves higher performance at lower temperatures.
-We perform ablation studies to analyze the impact of temperature on DISC.
-Typically, inference scaling methods achieve optimal performance at temperatures around 0.6–0.8, as increased temperature promotes sample diversity (Wang et al., 2024a).
-Surprisingly, however, DISC performs better at lower temperatures, as shown in Fig.
-11.
-This trend is in stark contrast to BoN (Fig.
-22), where higher temperatures are generally beneficial.
-We believe this phenomenon arises because DISC depends on accurately estimating the priority metric $h$ at each step.
-Lower temperatures reduce sample variance, leading to more reliable estimates of $h$, which in turn improves step selection.
-This is further supported by Fig.
-20, which shows that lower temperatures yield lower standard deviations per step, indicating increased sampling consistency.
-Additional details and analyses can be found in App.
-C.1.
-4.5 Self-Generated Validation Tests
-We also evaluate DISC in a more practical setting where a ground-truth reward model is unavailable for code generation  (Chen et al., 2022, 2023b; Zhou et al., 2024).
-Instead of relying on predefined test cases, we prompt the LLM to generate validation test cases based on the problem prompt.
-In real-world applications, manually curated ground-truth test cases are often costly to obtain, making self-generated validation a more scalable approach.
-The results, shown in Fig.
-12, indicate that DISC continues to scale better than other methods in this setting.
-Additional results and details are provided in App.
-D.3.
-Figure 12: Comparison of Pass@k performance on APPS with ground truth tests (right) and self generated validation tests (left) using gpt-4o-mini. DISC scales more effectively in both.
-4.6 Ablation on Base LLM Model
-We evaluate DISC across different LLMs, including open-source models.
-As shown in Fig.
-13 and Fig.
-27, DISC significantly enhances performance even for weaker models.
-Specifically, it improves Llama’s pass rate from 1% to 5.5%, a 550% relative increase, and Mistral’s from 0% to 3.5%, demonstrating substantial gains even from a nonzero baseline.
-Additional details and analyses are in App.
-C.3.
-Figure 13: Pass@k scaling curve for open-source models with DISC on APPS. DISC substantially improves performance across different LLMs, including Llama and Mistral.
-4.7 Interaction Between Priority Metric and DISC
-We conduct an ablation study to examine how the choice of priority metric affects DISC performance.
-In addition to the Q-based and Z-based priority metrics (DISC-Q and DISC-Z) introduced in Sec.
-3.2, we evaluate three baselines: DISC-R (random step selection), DISC-negQ, and DISC-negZ (which prioritize the opposite steps of DISC-Q and DISC-Z, respectively).
-As shown in Fig.
-14, the selection of a priority metric significantly impacts performance.
-Both DISC-Q and DISC-Z significantly outperform random selection and their inverse counterparts, demonstrating the effectiveness of their priority heuristics.
-Additional details and analysis are in App.
-C.2.
-Figure 14: Effect of priority metrics on DISC performance on APPS with gpt-4o-mini.
-Both Q-based and Z-based priority metrics outperform random selection and their inverses, highlighting their effectiveness.
-4.8 Ablation on Partition Fraction $\alpha$
-We conduct an ablation study to analyze the effect of the partition fraction $\alpha$ on DISC performance.
-As shown in Fig.
-15 and  31, the optimal range appears to be $0.15\leq\alpha\leq 0.25$.
-Lower partition fractions ($\alpha<0.5$) tend to perform better due to the asymmetric cost of sampling from different halves of the partition.
-Sampling from the first half requires generating more tokens, while the second half requires fewer, making it crucial to partition the first half more conservatively.
-Additional analysis are in App.
-C.4.
-Figure 15: Pass@k comparison of different DISC partition fractions $\alpha$ on APPS competition level. The range $0.15\leq\alpha\leq 0.25$ appears optimal.
-4.9 Search and DISC
-We demonstrate that search methods such as MCTS and beam search can be combined with DISC.
-As shown in Fig.
-47 in the Appendix, greedy search explores deeper partitions given the same search budget due to its greedy nature, while MCTS and beam search reach similar, shallower depths.
-However, MCTS allocates the search budget more effectively than beam search, leading to higher performance, as seen in Fig.
-16.
-Additional details and analysis are in App.
-E.
-Figure 16: Pass@k comparison of search methods combined with DISC on APPS with gpt-4o-mini. MCTS scales best, followed by greedy search, then beam search (with beam size 2).
-5 Related Work
-Inference scaling.
-Inference scaling has emerged as a dominant paradigm, driven by the introduction of o1- and r1-like chain-of-thought reasoning models (Snell et al., 2024; Brown et al., 2024; Manvi et al., 2024; Leea et al., 2025).
-Several works examine the trade-off between inference compute and training compute (Guan et al., 2025; Chen et al., 2024b).
-LLM inference often relies on decomposing complex problems into intermediate reasoning steps, as seen in chain-of-thought (CoT) prompting (Wei et al., 2022; Sprague et al., 2024; Wang & Zhou, 2024) and its variants (Kojima et al., 2022; Zhou et al., 2023; Wang et al., 2023b; Li et al., 2023).
-We extend inference scaling by introducing a new approach for adaptive compute allocation (Manvi et al., 2024).
-LLM reasoning and code generation.
-LLM reasoning and code generation are central tasks for inference scaling.
-Evolutionary inference scaling methods have been explored in program generation (Liventsev et al., 2023; Chen et al., 2023a; Romera-Paredes et al., 2024; Lehman et al., 2023; Hemberg et al., 2024).
-Domain-specific decomposition strategies have been applied in code generation, such as function-based decomposition (Chen et al., 2024a; Zenkner et al., 2024; Levin et al., 2025).
-More broadly, decomposition often involves prompting LLMs to generate subtask completions (Hernández-Gutiérrez et al., 2024; Khot et al., 2022; Dua et al., 2022), which differs from methods that refine a single LLM generation.
-Reinforcement learning and Monte Carlo methods.
-Unlike standard RL, our setting resembles a search problem where the goal is to identify the single highest-reward path.
-Cazenave (2009) demonstrated that nested Monte Carlo search can accelerate optimal pathfinding.
-Under the bandit setting, this can be formulated as identifying the arm with the highest maximum reward rather than the highest mean reward (Cicirello & Smith, 2005; Carpentier & Valko, 2014).
-6 Conclusion
-We introduce DISC, a dynamic decomposition framework that adaptively partitions solution steps based on difficulty, improving inference scaling by directing compute toward critical steps while balancing exploration and resource allocation.
-DISC seamlessly integrates with search-based methods such as MCTS and beam search, further enhancing performance.
-It also identifies challenging steps for LLMs, aiding curriculum learning, fine-tuning, and dataset augmentation.
-By dynamically adjusting partitioning based on available compute, DISC enables more adaptive and efficient reasoning in large language models, with broad implications for both training and inference optimization.
-7 Impact Statements
-This work on dynamic decomposition aims to advance the field of LLM by improving the efficiency and accuracy of large-scale inference methods, particularly in domains like code generation, mathematical problem-solving, and automated reasoning.
-By dynamically allocating computational resources to challenging reasoning steps, this approach has the potential to enhance AI-powered tools in software development, STEM education, and scientific research.
-While the core contributions of this paper focus on technical advancements, we acknowledge the broader societal implications of scaling inference techniques, including risks of over-reliance on AI in decision-making and the potential propagation of errors or biases.
-To mitigate these concerns, future work should emphasize robust validation, fairness, and human-in-the-loop deployment strategies, particularly in high-stakes applications.
-Nonetheless, no immediate or specific ethical concerns are identified in the context of this research.
-Appendix A Code implementation of DISC
-Python implementation of DISC def dynamic_decomposition(problem, model, reward_model, split_str, complete_solution, fraction, solution_budget, split_metric, stop_threshold=-float("inf"), stop_sum_score=1.0, stop_if_solved=False, ): """ ␣␣␣␣Decomposesthesolutionusingadynamicbinarysearchapproach ␣␣␣␣Args: ␣␣␣␣␣␣␣␣problem(Problem):Theproblemtosolve ␣␣␣␣␣␣␣␣model(Model):Themodeltouseforgeneration ␣␣␣␣␣␣␣␣reward_model(function):Therewardmodeltouseforscoring ␣␣␣␣␣␣␣␣split_str(function):Thefunctiontouseforsplittingastring ␣␣␣␣␣␣␣␣complete_solution(function):Thefunctiontouseforcompletingasolution ␣␣␣␣␣␣␣␣fraction(float):Thefractiontosplitthestring ␣␣␣␣␣␣␣␣solution_budget(int):Themaximumnumberofsolutionstogenerate ␣␣␣␣␣␣␣␣split_metric(function):Themetrictouseforsplitting ␣␣␣␣␣␣␣␣stop_threshold(float):Thethresholdtostopsplitting ␣␣␣␣␣␣␣␣stop_sum_score(float):Thesumscoretostopgeneratingcompletions ␣␣␣␣␣␣␣␣stop_if_solved(bool):Whethertostopiftheproblemissolved ␣␣␣␣""" # Initialize results and decomposition steps decomp_return = { "generated_solutions": [], "decomposition": [] } while len(decomp_return["generated_solutions"]) < solution_budget: # Combine all previous steps into an intermediate solution intermediate_solution = "".join([step["step_str"] for step in decomp_return["decomposition"]]) new_scores = [] best_solution = None best_completion = None best_score = -float("inf") sum_score = 0.0 # 1) Generate completions until we generate enough samples to estimate the split metric while sum_score < stop_sum_score: proposed_completion = complete_solution(problem, intermediate_solution, model) proposed_solution = intermediate_solution + proposed_completion decomp_return["generated_solutions"].append(proposed_solution) # Update scores proposed_score = reward_model(proposed_solution) new_scores.append(proposed_score) sum_score += proposed_score # Track the best solution if proposed_score > best_score: best_solution = proposed_solution best_score = proposed_score best_completion = proposed_completion # Stop early if problem is solved if stop_if_solved and proposed_score >= 1.0: decomp_return["decomposition"].append({"step_str": proposed_completion}) return decomp_return new_metric = split_metric(new_scores) last_metric = decomp_return["decomposition"][-1]["metric"] if decomp_return["decomposition"] else None # Determine the split target.
-We always split the step with the highest metric is_split_new_step = last_metric is None or new_metric >= last_metric split_target = decomp_return["decomposition"][-1]["step_str"] if not is_split_new_step else best_completion # 3) Attempt to split the target split_result = split_str(split_target, fraction) if not split_result: # If we can’t split the target, we’re done decomp_return["decomposition"].append({"step_str": best_completion, "metric": new_metric}) return decomp_return # Update decomposition based on split part1, part2 = split_result if is_split_new_step: decomp_return["decomposition"].append({"step_str": part1, "metric": new_metric}) # Stopping condition based on threshold if new_metric < stop_threshold: decomp_return["decomposition"].append({"step_str": part2}) return decomp_return else: decomp_return["decomposition"][-1] = {"step_str": part1, "metric": last_metric} return decomp_return
-Appendix B Pseudocode for DISC
-Algorithm 1 Dynamic Decomposition
-1:  Input: Problem instance ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}x}$, reward model ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r}$, partition function ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}f}$, LLM policy model ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\pi}$, partition fraction ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\alpha}$, solution budget ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$, priority metric ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}h}$, metric stopping precision ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$, sampling stopping threshold ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\sigma}$, is inference mode binference
-2:  Output: Final decomposition ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$
-[…]
-4:  # Decompose the solution recursively until we reach the desired precision ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$ or run out of budget ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$
-5:  while $|{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{generated\_solutions}|<{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$ do
-6:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}}\leftarrow\text{Concatenate}([\text{step.step\_str}\ \forall\ \text{step}\in{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}])$
-[…]
-8:     ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}}\leftarrow\varnothing$vartriangleright Record rewards of completions
-9:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{final}}}\leftarrow\text{None}$, ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\leftarrow\text{None}$, ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}\leftarrow-\infty$vartriangleright Track the best completion
-10:     # Step 1: Generate completions until we have enough samples to estimate the splitting metric. Here we use a geometric sampling distribution
-11:     while $\operatorname{sum}({\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}})<{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\sigma}$ do
-12:        ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\pi}(\cdot|{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}x},{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}})$
-13:        ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}}\oplus{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}$
-[…]
-15:        ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}\leftarrow{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r}({\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}})$
-16:        Append ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}$ to ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}}$
-17:        if ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}>{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}$ then
-18:           ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{final}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}}$, ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}$
-19:           ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}\leftarrow{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}$
-20:        end if
-21:        if $\textbf{b}_{\text{inference}}\ \textbf{and}\ {\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}=1.0$ then
-[…]
-23:           Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if problem is solved
-24:        end if
-25:     end while
-26:     # Step 2: Compute splitting metric
-27:     ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\leftarrow{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}h}({\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}})$
-28:     ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}[-1].{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z}\ \textbf{if}\ {\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}\neq\varnothing\ \textbf{else}\ -\infty$
-29:     # Step 3: Split the step with the higher metric
-30:     bsplit new step $\leftarrow$ ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\geq{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}$
-[…]
-32:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}},{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}f}({\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{target step}}},{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\alpha})$
-33:     if ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}}=\text{None}\ \textbf{or}\ {\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}=\text{None}$ then
-[…]
-35:        Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if we cannot do a finer split
-36:     end if
-37:     if $\textbf{b}_{\text{split new step}}$ then
-[…]
-39:        if ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}<{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$ then
-[…]
-41:           Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if all metrics are smaller than precision
-42:        end if
-43:     else
-44:        ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}[-1]\leftarrow\{\text{step\_str}:{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}},\text{metric}:{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}\}$vartriangleright Split last step
-45:     end if
-46:  end while
-47:  Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$
-Appendix C Ablation studies
-C.1 Ablation on Temperature
-We conduct an ablation study to analyze the effects of temperature on DISC and BoN.
-Temperature controls the randomness of token sampling in autoregressive models, influencing both exploration and consistency.
-Higher temperatures encourage more diverse outputs, whereas lower temperatures yield more deterministic generations.
-To examine its impact, we evaluate DISC and BoN on a 100-problem subset of APPS (the first 100 problems) using gpt-4o-mini.
-Figure 17 presents the Pass@token scaling curve for DISC across different temperatures.
-The results indicate that lower temperatures lead to improved performance, as DISC benefits from more deterministic step selection.
-Unlike BoN, which relies on broad solution sampling, DISC dynamically refines steps, making stable token probabilities advantageous.
-Figure 18 illustrates the frequency of actual partitions made by DISC at different temperatures.
-As temperature increases, the number of partitions fluctuates more, suggesting that high temperature introduces instability in step selection.
-Lower temperatures provide more structured decomposition, reducing unnecessary subdivisions.
-In Figure 19, we visualize the mean reward per step.
-The trend shows a linear increase in reward as step number grows, demonstrating that deeper decomposition results in progressively better solutions.
-This reinforces that DISC effectively allocates computation towards refining difficult steps.
-The mean standard deviation per step is shown in Figure 20.
-Lower temperatures yield lower standard deviations, confirming that DISC benefits from reduced variability in sample quality.
-This consistency allows for more reliable prioritization of difficult steps, enhancing overall inference efficiency.
-For comparison, Figure 22 and Figure 21 display Pass@token and Pass@k scaling curves for BoN across different temperatures.
-Unlike DISC, BoN achieves peak performance at a temperature around 0.6-0.8, balancing diversity and consistency.
-Higher temperatures increase exploration but degrade precision, while lower temperatures hinder sample diversity, reducing the probability of obtaining high-quality completions.
-These findings highlight the fundamental difference between DISC and BoN: DISC benefits from lower variance and stable decomposition, while BoN relies on broader exploration facilitated by moderate temperature settings.
-As a result, optimal temperature settings differ significantly between these methods, with DISC favoring deterministic sampling and BoN requiring a balance between diversity and coherence.
-Figure 17: Pass@token scaling curve for different temperatures on APPS using gpt-4o-mini. The lower the temperature, the stronger the DISC performance.
-Figure 18: Partition frequency of DISC with different temperatures on APPS using gpt-4o-mini
-Figure 19: Mean reward per step of DISC with different temperatures on APPS using gpt-4o-mini. The mean reward scales linearly with step number.
-Figure 20: Mean standard deviation per step of DISC with different temperatures on APPS using gpt-4o-mini. Lower temperature means lower average standard deviation.
-Figure 21: Pass@k scaling curve for different temperatures on APPS using gpt-4o-mini for BoN. A temperature around 0.6-0.8 leads to the best performance and balance between diversity and consistency.
-Figure 22: Pass@token scaling curve for different temperatures on APPS using gpt-4o-mini for BoN.
-A temperature around 0.6-0.8 leads to the best performance and balance between diversity and consistency.
-C.2 Ablation on Priority Metric $h$
-We analyze the effect of different priority metrics on DISC performance.
-We evaluate DISC using the first 200 competition-level APPS problems with gpt-4o-mini, setting the temperature to 0.8 for all experiments.
-The priority metric determines which steps are refined during recursive decomposition, impacting both efficiency and final solution quality.
-Figure 23 presents a token-level comparison of different priority metrics.
-Both DISC-Q and DISC-Z significantly outperform random selection and their inverse counterparts, demonstrating the importance of prioritizing high-value steps.
-Figure 24 illustrates the partition frequency under different priority metrics.
-We observe that effective metrics such as DISC-Q and DISC-Z lead to fewer, more meaningful partitions, whereas suboptimal strategies result in excessive, redundant partitioning.
-The relationship between mean reward and step number is shown in Figure 25.
-All tested metrics exhibit a strong correlation between increasing step depth and mean reward, indicating that decomposition progressively refines solutions.
-However, DISC-Q and DISC-Z achieve higher reward gains at earlier stages, suggesting that they prioritize the most impactful refinements.
-Finally, Figure 26 reports the standard deviation of rewards per step.
-Lower standard deviation suggests more stable solution quality, a property that DISC-Q and DISC-Z maintain better than random selection methods.
-This highlights their effectiveness in identifying and refining challenging steps efficiently.
-Overall, these results confirm that choosing an appropriate priority metric is crucial for DISC.
-While DISC-Q and DISC-Z consistently enhance inference efficiency and quality, random or inverse strategies lead to poorer performance due to misallocation of compute resources.
-Figure 23: Token level comparison of different priority metrics on DISC in the APPS setting with gpt-4o-mini. Both Q and Z based priority metrics perform well.
-Figure 24: Partition frequency of DISC with different priority metrics on APPS using gpt-4o-mini
-Figure 25: Mean reward per step of DISC with different priority metrics on APPS using gpt-4o-mini. All metrics display strong correlation between step depth and the mean reward.
-Figure 26: Mean standard deviation per step of DISC with different priority metrics on APPS using gpt-4o-mini. All metrics display correlation between step depth and the standard deviation.
-C.3 Model Ablation
-We investigate how different LLMs perform when used with DISC on 200 competition-level APPS problems, given a sample budget of 30.
-The groundtruth reward model was used to evaluate correctness, and all models were set to a temperature of 0.8.
-Due to the challenging nature of the benchmark, open-source models struggled to achieve strong performance independently.
-However, when paired with DISC, their performance significantly improved.
-Figure 27 presents the Pass@token scaling curve for open-source models using DISC.
-The results demonstrate that DISC substantially enhances the capabilities of these models, closing the gap between them and proprietary alternatives.
-Figure 28 visualizes the partition frequency of DISC with different open-source models.
-Compared to their standalone performance, the use of DISC led to more structured and effective decomposition, highlighting its adaptability to different architectures.
-The mean reward per step is shown in Figure 29.
-Similar to prior findings, we observe that deeper decomposition leads to increasingly higher rewards.
-Notably, even lower-capacity models benefit from DISC ’s ability to iteratively refine their solutions.
-Finally, Figure 30 presents the mean standard deviation per step. With DISC, the variance in performance is significantly reduced, resulting in more stable and reliable inference.
-Overall, these findings emphasize that DISC is a robust framework capable of enhancing inference performance across diverse LLMs, particularly those with limited standalone capabilities.
-Figure 27: Pass@token scaling curve for open source models with DISC on APPS. DISC also demonstrates strong performance gains with open source models.
-Figure 28: Partition frequency of DISC with open source models on APPS
-Figure 29: Mean reward per step of DISC with open source models on APPS
-Figure 30: Mean standard deviation per step of DISC with open source models on APPS
-C.4 Ablation on partition fraction $\alpha$
-We include some more analysis on the partition fraction here.
-Figure 31: Token level comparison of different DISC splitting fraction $\alpha$ on APPS competition level. $0.15\leq\alpha\leq 0.25$ seems to be optimal.
-Appendix D Experimental results extended
-D.1 APPS
-Figure 32: Sampling frequency of each step averaged over the problems on APPS with gpt-4o-mini. DISC seems to have a preferences for spending more compute on earlier found steps.
-D.2 Additional examples
-Below is another computed decomposition for the same problem as the one shown in the main text:
-Let $x$ be the length and $y$ be the width of the rectangle.
-Since the perimeter is 24 inches, we have $122x+2y=24\Rightarrow x+y=$ 12 .
-Therefore, we must maximize the area $xy$ subject to this constraint.
-We use the method of Lagrange multipliers. The Lagrangian is
-${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}\mathcal{L}(x,y,\lambda)=xy-\lambda(x+y-12).}$
-We differentiate with respect to $x,y,$ and $\lambda$ to obtain
-${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}\frac{\partial\mathcal{L}}{\partial x}=y-\lambda=0,\qquad\frac{\partial\mathcal{L}}{\partial y}=x-\lambda=0,\qquad\frac{\partial\mathcal{L}}{\partial\lambda}=x+y-12=0.}$
-We find that $x=y=\lambda$, so $x=y=\frac{12}{2}=6.$ Therefore, the maximum area of the rectangle is $6\cdot 6=\boxed{36}.$
-D.3 APPS with self-generated validation tests
-We examine DISC performance on APPS when using self-generated validation tests.
-All methods utilized the same set of self-generated validation tests to ensure fair comparisons.
-Each problem received 5-10 validation tests, with the exact number determined dynamically by the LLM.
-We evaluated a subset of 100 APPS problems, generating samples until the sample budget was exhausted or a correct solution was found.
-Figure 33 illustrates the Pass@token scaling curve, showing that DISC maintains strong scaling performance in this setting, though at a slightly lower rate compared to ground-truth verification.
-Figure 34 and Figure 35 compare actual and planned partition frequencies, respectively.
-The results indicate that DISC continues to make structured decompositions even with self-generated validation, preserving its efficiency.
-The mean reward per step, shown in Figure 36, follows a similar trend as in previous experiments, reinforcing that DISC effectively allocates compute resources for iterative refinement.
-Lastly, Figure 37 demonstrates that DISC maintains lower standard deviations in performance, indicating stable quality improvements across steps.
-Figure 33: Token level comparison of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
-DISC still scales better than other methods in this setting, albeit at a lower rate.
-Figure 34: Actual partition frequency of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
-Figure 35: Planned partition frequency of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
-Figure 36: Mean reward per step of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
-Figure 37: Mean standard deviation different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
-D.4 MATH500
-Completions for MATH500 include both the reasoning steps and the final answer.
-Since MATH500 contains more problems than APPS200 and MATH problems tend to be relatively easier, solution quality saturates quickly.
-Therefore, we use a lower sample budget of 10 for these experiments.
-Figure 38 presents the Pass@k performance for different decomposition methods on MATH500.
-We observe that all decomposition-based approaches achieve similar Pass@k performance, consistently outperforming BoN.
-This indicates that the structured nature of MATH problems allows multiple decomposition strategies to be effective.
-Despite similar Pass@k results, the true advantage of DISC lies in its token efficiency, as shown in Figure 6.
-DISC significantly reduces the number of tokens required to reach correct solutions compared to alternative methods, demonstrating its ability to allocate computational effort efficiently in mathematical reasoning tasks.
-Additionally, we analyze the partitioning behavior of DISC on MATH500.
-Figure 39 illustrates the actual partition frequency for different decomposition methods.
-The planned partitioning behavior, shown in Figure 40, further highlights how DISC effectively balances exploration and refinement.
-Finally, we present the mean standard deviation per step in Figure 41.
-Lower variance suggests that DISC produces more stable and reliable decompositions over multiple runs, reinforcing its robustness in both mathematical and program synthesis domains.
-Figure 38: Pass@k performance comparison for different decomposition methods on MATH500. DISC consistently outperforms BoN across different sampling budgets.
-Figure 39: Observed partition frequency of different decomposition methods on MATH500. DISC effectively segments problems into meaningful subcomponents.
-Figure 40: Planned partitioning strategy of different decomposition methods on MATH500. DISC ’s structured approach leads to more efficient problem breakdowns.
-Figure 41: Mean standard deviation per step for different decomposition methods on MATH500. Lower variance in DISC suggests more stable and reliable problem-solving steps.
-D.5 LiveCodeBench
-We evaluate DISC on LiveCodeBench, a benchmark designed for code generation tasks with a focus on real-world software development challenges.
-LiveCodeBench presents a unique set of problems requiring both reasoning and structured decomposition, making it a suitable testbed for evaluating DISC’s ability to refine and improve intermediate steps.
-Figure 42 shows the Pass@k comparison of different decomposition methods on LiveCodeBench.
-DISC consistently scales better than other decomposition methods, highlighting its ability to refine intermediate steps more effectively in complex coding scenarios.
-Figure 43 illustrates the observed partition frequency of different decomposition methods.
-The structured approach of DISC results in well-balanced decomposition across steps, reducing unnecessary partitioning while maintaining sufficient granularity for improved solution refinement.
-Figure 44 displays the planned partition frequency across methods.
-DISC dynamically determines the most effective partitions based on the evolving problem state, leading to more targeted and efficient decompositions.
-Finally, Figure 45 presents the mean standard deviation per step across decomposition methods.
-Lower variance in DISC suggests that it produces more stable and reliable decompositions, reinforcing its robustness for solving LiveCodeBench problems.
-Figure 42: Pass@k performance comparison for different decomposition methods on LiveCodeBench. DISC consistently outperforms other methods in structured problem refinement.
-Figure 43: Observed partition frequency of different decomposition methods on LiveCodeBench. DISC effectively balances problem segmentation while avoiding excessive partitioning.
-Figure 44: Planned partitioning strategy of different decomposition methods on LiveCodeBench. DISC dynamically adapts its partitioning to optimize search efficiency.
-Figure 45: Mean standard deviation per step for different decomposition methods on LiveCodeBench. Lower variance in DISC suggests more stable and reliable problem-solving steps.
-Appendix E Search and scaling
-E.1 Monte Carlo Tree Search (MCTS)
-Monte Carlo Tree Search (MCTS) is a widely used algorithm for sequential decision-making in large search spaces, particularly in applications such as game playing, planning, and inference scaling.
-The algorithm builds a search tree incrementally by simulating different sequences of actions and updating estimates of state quality.
-A key advantage of MCTS is its ability to balance exploration (discovering new states) and exploitation (refining promising ones) using a data-driven search process.
-The MCTS pipeline consists of four fundamental steps: selection, expansion, simulation, and backpropagation.
-E.1.1 Selection
-Starting from the root node representing the current state $\boldsymbol{s}$, MCTS iteratively traverses the search tree by selecting child nodes based on a selection policy.
-The most commonly used selection criterion is the Upper Confidence Bound for Trees (UCT), which balances exploration and exploitation:
-$UCT(\boldsymbol{s},\boldsymbol{d})=\hat{Q}(\boldsymbol{s},\boldsymbol{d})+c\sqrt{\frac{\ln\left(\sum_{\boldsymbol{b}}n(\boldsymbol{s},\boldsymbol{b})\right)}{n(\boldsymbol{s},\boldsymbol{d})}},$ (1)
-where $\hat{Q}(\boldsymbol{s},\boldsymbol{d})$ represents the estimated value of selecting action $\boldsymbol{d}$ from state $\boldsymbol{s}$, $n(\boldsymbol{s},\boldsymbol{d})$ is the visit count for this action, and $c$ is a hyperparameter controlling the trade-off between exploring new actions and favoring those with high past rewards.
-E.1.2 Expansion
-Once a leaf node (a previously unexplored state) is reached, the algorithm expands the tree by adding one or more new nodes.
-These new nodes represent potential future states $\boldsymbol{s}^{\prime}$ generated by sampling an action $\boldsymbol{d}$ from a predefined policy.
-This step broadens the search space and allows MCTS to evaluate new possibilities.
-E.1.3 Simulation
-Following expansion, the algorithm conducts a simulation (or rollout) from the newly added state.
-This step involves generating a sequence of actions according to a predefined policy until reaching a terminal state or an evaluation horizon.
-The outcome of the simulation, denoted as $v(\boldsymbol{s}^{\prime})$, provides an estimate of the quality of the new state.
-Depending on the application, this could represent a game result, an optimization score, or an inference accuracy metric.
-E.1.4 Backpropagation
-The final step involves propagating the results of the simulation back up the search tree to refine the estimated values of prior states and actions.
-Each node along the trajectory $\tau=[\boldsymbol{s}_{0},\boldsymbol{d}_{1},\boldsymbol{s}_{2},\dots,\boldsymbol{s}_{-1}]$ is updated iteratively:
-$\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t+1)}\leftarrow(1-\alpha_{n})\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t)}+\alpha_{n}\max\{\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t)},\hat{Q}(\boldsymbol{s}_{i+1},\boldsymbol{d}_{i+2})^{(t+1)}\},$ (2)
-where $\alpha_{n}$ is a learning rate that depends on the visit count, and the maximum function ensures that the best-performing trajectories are emphasized.
-MCTS has been widely adopted in inference scaling techniques due to its ability to efficiently allocate computational resources, focusing more on high-reward states while avoiding unnecessary exploration of unpromising regions.
-In later sections, we explore how MCTS can be combined with dynamic decomposition to further optimize inference scaling.
-E.1.5 Combining Dynamic Decomposition with MCTS
-MCTS can be enhanced by integrating dynamic decomposition, where each node in the search tree represents a decomposition of the problem into steps.
-Instead of treating states as atomic decisions, we recursively decompose reasoning steps, dynamically adjusting granularity based on difficulty.
-In this framework:
-  • Each node in the MCTS tree represents a partial decomposition of the problem, with child nodes corresponding to alternative step partitions.
-  • Branching occurs by generating candidate next steps using dynamic decomposition, allowing finer steps for complex regions while maintaining efficiency for simpler ones.
-  • The selection step prioritizes nodes that represent more promising decompositions, dynamically refining challenging areas through recursive subdivision.
-  • The backpropagation step ensures that decompositions leading to high-quality solutions are reinforced, helping the search tree converge toward optimal inference paths.
-By integrating dynamic decomposition with MCTS, we efficiently allocate compute to the most critical reasoning steps, improving inference quality while maintaining computational efficiency.
-E.2 Beam Search
-Beam search is a heuristic search algorithm commonly used in inference tasks where computational efficiency is a priority.
-Unlike exhaustive search methods, beam search maintains only the top $k$ best candidates at each step, making it an effective strategy for structured prediction problems and sequential decision-making.
-At each iteration:
-  • The algorithm selects the $k$ most promising partitions from the previous step based on an evaluation metric.
-  • Each selected partition is expanded by generating possible next-step samples.
-  • The newly generated partitions are ranked, and only the top $k$ candidates are retained for the next iteration.
-  • This process continues until a stopping criterion is met, such as reaching a predefined depth or finding a sufficiently high-quality solution.
-Beam search provides a computationally efficient way to explore structured solution spaces while maintaining high-quality search trajectories.
-By integrating beam search with dynamic decomposition, we ensure that inference computation is allocated efficiently, focusing on the most promising reasoning paths at each step.
-E.3 Additional Results and Analysis
-Experiments comparing different search methods were conducted on a 100-problem subset of the APPS dataset (first 100 problems) using GPT-4o-mini.
-All methods used a temperature of 0.2, with $\alpha=0.15$, Q priority metric, and $\sigma=1.0$.
-Token-level comparison: As shown in Figure 46, MCTS scales best among the tested methods, demonstrating superior efficiency in identifying promising partitions.
-Greedy search follows closely, while beam search exhibits the slowest scaling.
-Partition frequency analysis: Figure 47 reveals that greedy search explores to greater depths within the same sampling budget.
-This suggests that greedy search prioritizes deep refinements, whereas MCTS and beam search balance depth with breadth.
-Step variance analysis: Figure 48 illustrates that all search methods display decreasing standard deviation with increasing search depth.
-This trend indicates that deeper searches converge towards stable, high-quality partitions, reinforcing the benefits of dynamic decomposition.
-These results highlight the trade-offs between search methods: MCTS offers robust exploration-exploitation balance, greedy search favors depth-first refinement, and beam search provides a structured yet computationally constrained approach.
-The integration of dynamic decomposition further enhances these search strategies by adaptively allocating computational resources to critical reasoning steps.
-Figure 46: Token level comparison of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. MCTS scales best, followed by greedy search, followed by beam search.
-Figure 47: Actual partition frequency of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. Greedy is able to search to higher depths given the same sampling budget.
-Figure 48: Mean standard deviation of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. All search methods display decreasing standard deviation with search depth.
+Appendix H Compute Resources Used
+All evaluations involving OpenAI proprietary models (e.g., GPT-4o-mini, GPT-3.5) were performed via the OpenAI API.
+These API calls were made from standard desktop machines using CPU-only inference and incur no dependency on local GPU availability, making the method broadly accessible for replication.
+For open-source model experiments (e.g., LLaMA-3.1-8B-Instruct, Mistral, Qwen), inference was conducted on a single NVIDIA A100 GPU.
+All such experiments were executed sequentially on this GPU, which has 80GB of memory, ensuring consistent and reproducible runtime characteristics across model families.
+This setup reflects the lightweight computational overhead of our method and demonstrates that DISC can scale effectively even in constrained or CPU-only environments when using hosted APIs.
+Appendix I Impact Statement
+Positive Societal Impacts
+This work introduces a general and lightweight method—Dynamic Decomposition (DISC)—that significantly improves inference efficiency for large language models (LLMs) without requiring additional training, domain-specific engineering, or specialized hardware.
+By enabling better performance using fewer samples and tokens, DISC reduces the cost and environmental footprint associated with LLM deployment, making advanced reasoning models more accessible to research groups and developers with limited compute budgets.
+Additionally, DISC’s ability to prioritize critical reasoning steps has the potential to improve the transparency and interpretability of LLM outputs, which is beneficial in high-stakes applications such as education, scientific reasoning, and assistive tools for programming or mathematics.
+Its plug-and-play compatibility with open-source models also democratizes access to high-performance inference techniques.
+Negative Societal Impacts
+As with all improvements in LLM inference capabilities, this work may accelerate the deployment of models in settings where societal risks are not fully mitigated—such as the automated generation of persuasive misinformation, cheating in educational contexts, or manipulation via high-fidelity language generation.
+Furthermore, by making LLM inference more efficient, DISC could contribute to increased usage of models without corresponding increases in ethical oversight or alignment safeguards.
+Careful integration of this method should therefore include responsible use policies, limitations on deployment domains, and alignment with values such as transparency, fairness, and accountability.